### v3

# convergence.bib

Abstract
We study a discrete and continuous version of the spectral Dirichlet problem in an open bounded connected set $\Omega\subset\mathbb{R}^{d}$, in dimension $d\geq 2$. More precisely, consider the simple random walk on $\mathbb{Z}^{d}$ killed upon exiting the (large) bounded domain $\Omega_{N}=(N\Omega)\cap\mathbb{Z}^{d}$. We let $P_{N}$ its transition matrix and we study the properties of its ($L^{2}$-normalized) principal eigenvector $\phi_{N}$, also known as ground state. Under mild assumptions on $\Omega$, we give regularity estimates on $\phi_{N}$, namely on its $k$-th order differences, with a uniform control inside $\Omega_{N}$. We provide a completely probabilistic proof of these estimates: our starting point is a Feynman–Kac representation of $\phi_{N}$, combined with gambler’s ruin estimates and a new “multi-mirror” coupling, which may be of independent interest. We also obtain the same type of estimates for the first eigenfunction $\varphi_{1}$ of the corresponding continuous spectral Dirichlet problem, in relation with a Brownian motion killed upon exiting $\Omega$. Finally, we take the opportunity to review (and slightly extend) some of the literature on the $L^{2}$ and uniform convergence of $\phi_{N}$ to $\varphi_{1}$ in Lipschitz bounded domains of $\mathbb{R}^{d}$, which can be derived thanks to our estimates.
Keywords: potential theory, spectral Dirichlet problem, finite difference method, random walk, mirror coupling, gambler’s ruin, Harnack inequalities.
2020 Mathematics subject classification: 60G50, 65L12
1 Introduction
The aim of this paper is to revisit a classical spectral Dirichlet eigenvalue problem from a probabilistic perspective. Specifically, we examine a discrete version of the problem that naturally arises when studying random walks that are killed upon exiting a large domain. Our primary goal is to provide probabilistic proofs (primarily based on gambler’s ruin estimates and coupling techniques) of regularity properties for the principal eigenfunction both at the discrete and at the continuous level — in the spirit of difference estimates and Harnack inequalities, see [lawlerRandomWalkModern2010, §6.3]. Most notably, our proofs is based on a Feynman–Kac representation for the principal eigenfunction (see e.g. [dynkin2002diffusions, §6.3]), combined with a novel “multi-mirror” coupling to derive estimates on its $k$-th order differences.
The use of coupling arguments has already proven to be very helpful in the context of spectral theory, see [atarburdzy2004, kendall1989] for the study of Neumann eigenfunctions, see also [burdzykendall2000, chen1998] for some overviews of other possible applications. Let us quote here Kendall [kendall1989] to highlight the interest of probabilistic techniques for spectral problems: “Probability is useful here because it enables a coupling argument, working on individual sample paths. Analytical arguments via the heat equation tend to integrate the manifold variety of paths from point to point, and thus lose flexibility.” To our knowledge, such techniques have not been used for Dirichlet eigenfunctions and their discrete approximations, and we aim at applying Kendall’s philosophy in this context.
Let us mention that the present paper initiated because we could not find a proper reference for the uniform convergence of the discrete eigenvector to its continuous counterpart; we derive this convergence as a corollary of our estimates. This is a classical topic in numerical analysis and potential theory, but its literature is extensive and can be challenging to navigate for non-experts. To help in this regard, we therefore take the opportunity to review (and complete) some of the literature, see Section 2.3 below.
1.1 The Dirichlet eigenvalue problem and a discrete approximation
Spectral Dirichlet problem.
Let $\Omega$ be an open, bounded and connected set of $\mathbb{R}^{d},d\geq 2$; we assume for simplicity that it contains the origin. Consider the spectral Dirichlet problem on $\Omega$, of finding $\mu\in\mathbb{R}\,,\;v\in L^{2}(\Omega)$ such that
$\begin{cases}-\Delta v=\mu\,v&\quad\text{ on }\Omega\,,\\ \quad\ \ v=0&\quad\text{ on }\partial\Omega\,,\end{cases}$ (1.1)
with $\Delta=\displaystyle\sum_{i=1}^{d}\frac{\partial^{2}}{\partial x_{i}^{2}}$ the usual Laplace operator.
By spectral theory it is classical that the spectrum of the Dirichlet Laplacian is discrete with positive eigenvalues; we refer for instance to [henrot2006extremum, Thm. 1.2.5]. We denote by $(\mu_{k},\varphi_{k})_{k\geq 1}$ the ordered eigenvalues and $L^{2}$-normalized eigenfunctions. Additionally, the smallest eigenvalue $\mu_{1}$ is simple with associated eigenfunction $\varphi_{1}$ of constant sign, see e.g. [henrot2006extremum, Sec. 1.3.3] or more plainly [Borthwick2021, Thm. 6.34].
We will mostly consider the case of a Lipschitz domain $\Omega$, i.e.  such that its boundary $\partial\Omega$ is locally the graph of a Lipshitz function, i.e. $x_{d}=\psi(x_{1},\ldots,x_{d-1})$ (in appropriate local coordinates) for some Lipschitz function $\psi$. This includes for instance smooth domains, with $\mathcal{C}^{1}$ boundary, it also allows for (reentrant) corners in the boundary, but not for cusp. This Lipschitz domain condition is classical in the potential theory literature, and ensures for instance that there is a bounded trace operator $\gamma:H^{1}(\Omega)\to H^{1/2}(\partial\Omega)$, see [mclean2000strongly, Th. 3.38]; in other words, we can give a pointwise meaning to $\varphi_{k}=0$ on $\partial\Omega$ (see our estimates or Proposition 2.1). Let us point to [agranovich2013spectral] for a review of spectral problems in domains that are Lipschitz, see in particular its Section 4 for the spectral Dirichlet eigenvalue problem.
A discrete version of the problem.
One can discretize the Dirichlet eigenvalue problem (1.1) in several ways, and we focus on the finite-difference method, because of its clear relation with the simple random walk (see Section 5 below).
Let $h>0$ be a mesh size and consider $\mathbb{Z}_{h}^{d}\vcentcolon=(h\mathbb{Z}^{d})$. We then define the discretized versions of $\Omega$ and $\partial\Omega$: we let $\Omega^{(h)}$ be the connected component of 00 in $\Omega\cap\mathbb{Z}_{h}^{d}$ (viewed as a subset of the lattice $\mathbb{Z}_{h}^{d}$) and $\partial\Omega^{(h)}\vcentcolon=\{x\in h\mathbb{Z}^{d}\setminus\Omega^{(h)}\,,|x-y|<h\}$. Let also $\Delta^{(h)}$ be the discrete analogue of the Laplacian $\Delta$: for $v:\mathbb{Z}_{h}^{d}\longrightarrow\mathbb{R}$,
$\Delta^{(h)}v(x)=h^{-2}\sum_{e\in\{\pm e_{i},1\leq i\leq d\}}\big{(}v(x+he)-v(x)\big{)}\,,$
where $e_{i}$ is the $i$-th vector of the canonical base of $\mathbb{Z}^{d}$. Discrete potential theory results in Lipschitz domains exist, and let us give a few examples: for instance, [varopoulos2001potential] provides kernel and Green function estimates (in particular near the boundary); [varopoulos2009discrete] gives some (optimal) $L^{\infty}$ convergence for solution of the classical Dirichlet problem (i.e. $\Delta u=0$ in $\Omega$, with $u=f$ on the boundary); [MS19] investigates the Martin boundary of unbounded globally Lipschitz domains.
Here, we consider the discrete analogue of the spectral Dirichlet problem (1.1):
$\begin{cases}-\Delta^{(h)}v^{(h)}=\mu^{(h)}\,v^{(h)}&\quad\text{ on }\Omega^{(h)}\,,\\ \qquad\ \ v^{(h)}=0&\quad\text{ on }\partial\Omega^{(h)}\,,\end{cases}$ (1.2)
with $v^{(h)}:\Omega^{(h)}\cup\partial\Omega^{(h)}\to\mathbb{R}$. We denote by $(\mu_{k}^{(h)},\varphi_{k}^{(h)})_{k\geq 1}$ the ordered eigenvalues and $L^{2}$-normalized eigenfunctions of (1.2) (see Remark 1.1 below), and it is here easy to see (by the Perron–Frobenius theorem) that the first eigenvalue $\mu_{1}^{(h)}$ is simple with an associated eigenfunction $\varphi_{1}^{(h)}$ which is positive in $\Omega^{(h)}$.
Then the finite difference method shows that, for “smooth” domains $\Omega$, for any $k\geq 1$, $\lim_{h\downarrow 0}\mu_{k}^{(h)}=\mu_{k}$ and $\lim_{h\downarrow 0}\varphi_{k}^{(h)}=\varphi_{k}$ in $L^{2}$ and in the sup norm, with explicit rates. We refer to Section 2.3 below for a more detailed discussion.
Remark 1.1 (About the normalization).
In this paper, we consider the $L^{2}$-normalized eigenfunctions, i.e. such that
$\|\varphi_{k}\|_{L^{2}}^{2}\vcentcolon=\int_{\Omega}\varphi_{k}(x)^{2}\mathrm{d}x=1\quad\text{ and }\quad\|\varphi_{k}^{(h)}\|_{L^{2},h}^{2}\vcentcolon=h^{d}\sum_{x\in\Omega^{(h)}}\varphi_{k}^{(h)}(x)^{2}=1\,.$
One could also study $L^{1}$-normalized eigenfunctions $\phi_{k}$, $\phi_{k}^{(h)}$, i.e. such that $\int_{\Omega}|\phi_{k}(x)|\mathrm{d}x=1$ and $h^{d}\sum_{x\in\Omega^{(h)}}|\phi_{k}^{(h)}(x)|=1$, or some $L^{\infty}$ or point-normalization, setting for instance the value at 00 to be equal t  $1$. All these normalization turn out to be equivalent in our context, so we will focus later on the $L^{2}$-normalized version of the eigenfunction, see Remark 2.1 below.
Our main results in a nutshell.
In the rest of the paper, we focus on the discrete principal eigenvalue $\mu_{1}^{(h)}$ and eigenfunction $\varphi_{1}^{(h)}$, and on their continuous counterpart. Our goal is to otbain, via probabilistic techniques, the following estimates.
Theorem 1.1.
Suppose that $\Omega$ is a Lipschitz domain (see Assumption 2 below for a precise definition). Then, there exist some (explicit) $p\in(0,1]$ and some constant $C>0$ such that, for all $k\geq 0$:
$\begin{split}\Big{|}\frac{\partial^{k}}{\partial x_{i_{1}}\cdots\partial x_{i_{k}}}\varphi_{1}(x)\Big{|}\leq(Ck)^{k}d(x,\partial\Omega)^{p-k}\,,&\qquad\forall\,x\in\Omega\,,\\ \Big{|}D_{i_{1},\cdots i_{k}}\varphi_{1}^{(h)}(x)\Big{|}\leq(Ck)^{k}d(x,\partial\Omega)^{p-k}\,,&\qquad\forall\,x\in\Omega^{(h)}\text{ with }d(x,\partial\Omega)>2kh\,,\end{split}$
where $D_{i_{1},\ldots,i_{k}}$ is the $k$-th order finite difference operator in direction $(i_{1},\ldots,i_{k})$, see its definition (2.6) below. Note that for $k=0$, the results are simply about the eigenfunctions $\varphi_{1},\varphi_{1}^{(h)}$.
These derivative or difference estimates should be classical results, but we were not able to find a reference in the literature, especially in the case of Lipschitz domains. Let us simply mention that a notable feature of the bounds is that they control the eigenfunction and their derivatives near the boundary $\partial\Omega$. Comparing the discrete and continuous estimates, loosely speaking the results can be interpreted as telling that $\varphi_{1}^{(h)}$ has the same regularity as its continuous counterpart $\varphi_{1}$, uniformly in $h\in(0,1)$. In the following, we split the presentation of Theorem 1.1 into several parts, in particular to highlight how the parameter $p$ depends on the domain $\Omega$.
1.2 Simple random walk and Brownian motion in a (large) bounded domain
Our proofs rely on coupling techniques, so let us introduce briefly the probabilistic objects that we use, and their relation to the Dirichlet spectral problems (1.1)-(1.2).
Simple random walk in $\Omega_{N}$.
The discretized eigenvalue problem (1.2) appears naturally in the context of random walks on $\mathbb{Z}^{d}$, as follows. Let $N\geq 1$ be a large integer and consider the domain $\Omega_{N}\subset\mathbb{Z}^{d}$ defined as the connected component of 00 in $(N\Omega)\cap\mathbb{Z}^{d}$. In the case where the boundary $\partial\Omega$ is $\mathcal{C}^{1}$, then one can simply define
$\Omega_{N}=(N\Omega)\cap\mathbb{Z}^{d}\,,$
at least for $N$ large enough. We also let $\partial\Omega_{N}\vcentcolon=\{x\in\mathbb{Z}^{d}\setminus\Omega_{N},\,\exists y\in\Omega_{N}\text{ with }y\sim x\}$. One can simply interpret $\Omega_{N}$ as $h^{-1}\Omega^{(h)}$ with $h=\frac{1}{N}$, but in the context of random walks we prefer to work with $\Omega_{N}\subset\mathbb{Z}^{d}$ rather than $\Omega^{(h)}$.
We consider $\mathbb{Z}^{d}$ endowed with the Euclidean norm ||$\cdot$|| ⋅ | and we denote by $d(\cdot,\cdot)$ its graph distance (given by the $1$-norm $|\cdot|_{1}$); we denote $y\sim x$ if $x,y$ are nearest-neighbors, i.e. $d(x,y)=1$. We also denote by $\Delta_{d}$ the graph Laplacian on $\mathbb{Z}^{d}$, i.e. $\Delta_{d}v(x)=\sum_{y\sim x}(v(y)-v(x))$. Consider the matrix $P_{N}$ of the (nearest-neighbor) simple random walk on $\mathbb{Z}^{d}$ killed upon exiting $\Omega_{N}$, namely
$P_{N}(x,y)=\begin{cases}\displaystyle\frac{1}{2d}\quad&\text{ for }x\sim y,\,x,y\in\Omega_{N},\\ \displaystyle 0&\text{ otherwise}.\end{cases}$
We then focus on the principal eigenvalue $\lambda_{N}=\lambda_{\Omega_{N}}$ of $P_{N}$ and its associated $L^{2}$-normalized eigenvector $\phi_{N}=\phi_{\Omega_{N}}$ (which is positive on $\Omega_{N}$),
$P_{N}\phi_{N}=\lambda_{N}\phi_{N}\,,\qquad\text{ with }\ \|\phi_{N}\|_{L^{2},\frac{1}{N}}^{2}\vcentcolon=\frac{1}{N^{d}}\sum_{x\in\Omega_{N}}\phi_{N}(x)^{2}=1\,.$ (1.3)
By definition of $P_{N}$, one easily verifies that $I_{d}-P_{N}=-\frac{1}{2d}\Delta_{d}$, so $\lambda_{N},\phi_{N}$ are related to (1.2) with $h\vcentcolon=\frac{1}{N}$ in the following way: $\lambda_{N}\vcentcolon=1-\frac{1}{2d}\mu_{1}^{(h)}$ and $\phi_{N}(x)=\varphi_{1}^{(h)}(hx)$.
One of the goal of the paper is to study properties of $\phi_{N}$ which can prove useful when considering the random walk conditioned to remain forever inside $\Omega_{N}$. Indeed, one can introduce the Doob’s $\phi_{N}$-transform of the simple random walk, defined by the transition kernel:
$\tilde{p}_{N}(x,y)\vcentcolon=\lambda_{N}^{-1}\frac{\phi_{N}(y)}{\phi_{N}(x)}\;\frac{1}{2d}\mathbbm{1}_{\left\{x\sim y\right\}}\qquad\text{ for }x,y\in\Omega_{N}\,.$ (1.4)
By standard Markov chain theory (see e.g. [lawlerRandomWalkModern2010, App. A.4.1]), the transition kernel (1.4) is the limit, as $t\to+\infty$, of the transition kernels of the SRW conditioned to stay in $\Omega_{N}$ until time $t$ (see also (3.2) below). We thus refer to the Markov process with transition kernel (1.4) as the random walk confined in $\Omega_{N}$. This is also referred to as the quasi-stationary distribution (QSD), see [collet2013quasi] for an overview. Hence, obtaining detailed estimates on $\phi_{N}$ gives precious information on the transition kernel $\tilde{p}_{N}$.
Notice also that (1.4) can also be interpreted as the transition kernel of a random walk among conductances $c_{N}(x,y)\vcentcolon=\phi_{N}(x)\phi_{N}(y)\mathbbm{1}_{\{x\sim y\}}$, i.e. we can rewrite
$\tilde{p}_{N}(x,y)=\frac{c_{N}(x,y)}{\sum_{z\sim x}c_{N}(x,z)}\qquad\text{ for}x,y\in\Omega_{N}\,.$
Let us also note the stationary distribution of the random walk confined in $\Omega_{N}$ is given by $\pi_{N}(x)=\phi_{N}^{2}(x)$, the $L^{2}$-normalization of $\phi_{N}$ making $\pi_{N}$ a probability distribution. We refer to [lawlerRandomWalkModern2010, App. A.4.1] for some details.
Brownian motion in $\Omega$.
Analogously, one can also consider the Brownian motion killed upon exiting $\Omega$ and the Brownian motion conditioned to remain forever inside $\Omega$, see [pinsky1985convergence] or [chung2012brownian, Ch. 2 and Ch. 5]. Let us briefly present its relation with the spectral Dirichlet problem (1.1).is
A Brownian motion killed when exiting $\Omega$ is a Markov process with transition kernel $\rho_{t}^{\Omega}(x,A)=\mathbf{P}_{x}(X_{t}\in A,H_{\partial\Omega}>t)$, where $(X_{s})_{s\geq 0}$ is a standard Brownian motion (starting from $x$ under $\mathbf{P}_{x}$) and $H_{\partial\Omega}=\inf\{s>0,X_{s}\in\partial\Omega\}$ is the exit time of $\Omega$. Its generator is the Dirichlet Laplacian on $\Omega$, i.e. with boundary condition considered in (1.1).
Analogously to (1.4), the Brownian motion conditioned to remain forever inside $\Omega$ (or QSD) is the Doob’s $\varphi_{1}$-transform of the killed Brownian motion, where $\varphi_{1}$ is the first eigenvalue of the Dirichlet Laplacian, see (1.1). In other words, this is the Markov process with transition density
$\tilde{\rho}_{t}(x,y)=\frac{\varphi_{1}(y)}{\varphi_{1}(x)}e^{\mu_{1}t}\rho_{t}^{\Omega}(x,y)\,,$ (1.5)
where $\rho_{t}^{\Omega}(x,y)$ is the density transition of the killed Brownian motion.
2 Main results: regularity properties of the eigenfunctions
The main results of this paper give some properties of the first $L^{2}$-normalized discrete eigenvector $\phi_{N}$, under some mild condition on the domain $\Omega$; they have corresponding results for the continuous eigenfunction $\varphi_{1}$. We will work with two different assumptions on the regularity of the boundary of $\Omega$, which affects the regularity of eigenfunctions near $\partial\Omega$.
Main assumptions on the domain.
A first (stronger) assumption that we will use is that the set $\Omega$ has a positive reach, which informally tells that one can roll a ball on the outer boundary of $\Omega$. This includes the case of $\mathcal{C}^{2}$ boundaries, but also allows for non-reentrant corners.
Assumption 1 (Positive reach).
There exists some $\varepsilon>0$ such that for any $x\in\partial\Omega$ there is some $z$ with $d(z,x)=\varepsilon$ such that $B(z,\varepsilon)\cap\Omega=\emptyset$.
Our second, weaker, condition is a “uniform exterior cone” condition; this is a standard assumption in the literature, corresponding to a uniform version on Poincaré’s cone property. For bounded domains, this is actually equivalent to considering a Lipschitz domain, see [grisvard2011elliptic, Thm. 1.2.2.2], but we prefer the exterior cone formulation since it appears naturally in our proofs (it has a probabilistic interpretation, see for instance in [deblassie1987]).
[…]
$\alpha_{r}(x)\vcentcolon=\sup\big{\{}\alpha\,:\,\exists y\in\mathbb{S}^{d-1}\text{ such that }\mathcal{C}_{\alpha,y}(y)\cap B(x,r)\cap\Omega=\emptyset\big{\}}\,.$ (2.1)
We can now state our exterior cone condition.
Assumption 2 (Uniform exterior cone condition).
There is some radius $r=r_{\Omega}>0$ such that
$\alpha=\inf_{x\in\partial\Omega}\alpha_{r}(x)>0\,.$ (2.2)
In other word, there is an $\alpha=\alpha_{\Omega}>0$ such that, for all $x\in\Omega$, there exists an exterior open cone $\mathcal{C}_{\alpha}(x)$ with vertex $x$ and angle $\alpha$ such that $\mathcal{C}_{\alpha}(x)\cap B(x,r)\cap\Omega=\emptyset$.
These two assumptions are related to the regularity properties of the eigenfunctions, and will in particular impact how easy it is for a random walk or a Brownian motion to avoid exiting the domain $\Omega_{N}$ when near the boundary. The proofs and the results are somehow cleaner under Assumption 1, so we start our presentation in this case. We show along our paper how our results are affected assuming Assumption 2 instead of Assumption 1.
2.1 Under the positive reach condition, Assumption 1
Let us first start with some bound on $\phi_{N}(x)$ that depends on the distance to the boundary. We mention that our proof relies on simple gambler’s ruin estimates (i.e. probability to avoid a ball).
Proposition 2.1.
Suppose that $\Omega$ satisfies Assumption 1. Then there is a constant $C>0$ such that
$0<\phi_{N}(x)\leq C\,\frac{d(x,\partial\Omega_{N})}{N}\,,\qquad\forall\,x\in\Omega_{N}\,.$ (2.3)
In particular, we have $\sup_{x\in\Omega_{N}}|\phi_{N}(x)|\leq C\,\mathrm{diam}(\Omega)$.
Note that, in the context of the eigenvalue problem in the form (1.2), this proposition gives Theorem 1.1 for $k=0$.
Remark 2.1.
Proposition 2.1 actually shows that the sup-norm of $\phi_{N}$ is controlled by its $L^{2}$-norm (equal to $1$), uniformly in $N\geq 1$. Since the $L^{2}$-norm is obviously controlled by the sup-norm, we obtain that if $\Phi_{N}$ is the $L^{\infty}$-normalized principal eigenfunction, then we have that $\Phi_{N}=C_{N}\phi_{N}$ with a constant $C_{N}$ uniformly bounded away from 00 and $\infty$. Similar comparisons hold for the $L^{1}$ norm.
Let us now state our main results, that shows that $\phi_{N}$ varies regularly inside $\Omega_{N}$; our proofs rely on a Feynman–Kac representation, together with a probabilistic, “multi-mirror”, coupling argument. For pedagogical purposes, let us already give a first control on differences of $\phi_{N}$, and turn afterwards to higher-order differences (we need to introduce further notation); our proof follows the same pedagogical scheme, starting with a first-order difference estimate where the mirror coupling is simple.
Theorem 2.2 (Differences of $\phi_{N}$).
Suppose that $\Omega$ satisfies Assumption 1. Then there is a constant $C>0$ (that depends only on the domain $\Omega$) such that
$\big{|}\phi_{N}(x)-\phi_{N}(y)\big{|}\leq C\,\frac{d(x,y)}{N}\,,\qquad\forall\,x,y\in\Omega_{N}\,.$ (2.4)
Let us now define the $k$-th order difference of a function $\psi:\mathbb{Z}^{d}\to\mathbb{R}$. For $i\in\{1,\ldots,d\}$, let $D_{i}$ be the difference operator in direction $i$, defined as
$D_{i}\psi(x)=\frac{1}{2}\big{(}\psi(x+e_{i})-\psi(x-e_{i})\big{)}\,,$ (2.5)
with $e_{i}$ the $i$-th vector of the canonical basis. (For instance, note that Theorem 2.2 shows that $|D_{i}\phi_{N}(x)|\leq C/N$ uniformly in $x\in D_{N}$.) Then, for $k\geq 2$, we define the $k$-th order difference in directions $i_{1},\ldots,i_{k}\in\{1,\ldots,d\}$ by setting
$D_{i_{1},\ldots,i_{k}}=D_{i_{1}}\cdots D_{i_{k}}\,.$
In fact, a simple iterative argument shows that, for any $i_{1},\ldots,i_{k}\in\{1,\ldots,d\}$,
$D_{i_{1},\ldots,i_{k}}\psi(x)=D_{i_{1}}\cdots D_{i_{k}}\psi(x)=\frac{1}{2^{k}}\sum_{\alpha\in\{+1,-1\}^{k}}\mathrm{sign}(\alpha)\;\psi\Big{(}x+\sum_{j=1}^{k}\alpha_{j}e_{i_{j}}\Big{)}\,,$ (2.6)
where $\mathrm{sign}(\alpha)=(-1)^{m}$ with $m$ the number of “$-1$” in $\alpha$.
Remark 2.2.
Let us stress that the $k$-th order derivatives are finite-difference schemes for the partial derivatives. Indeed, for a function $\psi:\mathbb{R}^{d}\to\mathbb{R}$ and $h>0$, we can define
$D^{(h)}_{i_{1},\ldots,i_{k}}\psi\vcentcolon=D_{i_{1},\ldots,i_{k}}\psi_{h}\qquad\text{ with }\ \psi_{h}:\begin{array}[]{cll}\mathbb{Z}^{d}&\to\mathbb{R}^{d}\\ x&\mapsto h^{-1}\psi(hx)\end{array}\,.$ (2.7)
Then, if $\psi$ is $k$ times differentiable at $x$, we have that $\frac{\partial^{k}}{\partial x_{i_{1}}\cdots\partial x_{i_{k}}}\psi(x)=\lim_{h\downarrow 0}D^{(h)}_{i_{1},\ldots,i_{k}}\psi(x)$.
We then have the following result on the higher-order differences of $\phi_{N}$ (for $k=1$ it recovers Theorem 2.2, but we state it separately for readability), which is the discrete part of Theorem 1.1.
Theorem 2.3 ($k$-th order differences of $\phi_{N}$).
Suppose that $\Omega$ satisfies Assumption 1. Then there is a constant $C>0$ (that depends only on the domain $\Omega$) such that, for any $k\geq 1$ and any $i_{1},\ldots,i_{k}\in\{1,\ldots,d\}$, we have
$\big{|}D_{i_{1},\ldots,i_{k}}\phi_{N}(x)\big{|}\leq\frac{(Ck)^{k}}{N^{k}}\Big{(}\frac{d(x,\partial\Omega_{N})}{N}\Big{)}^{1-k}\,,\qquad\forall\,x\in\Omega_{N}\,.$ (2.8)
Remark 2.3.
We could also treat directional $k$-th order differences, defining $D_{i^{+}}$, $D_{i^{-}}$ by $D_{i^{+}}\psi(x)=\psi(x+e_{i})-\psi(x)$ and $D_{i^{-}}\psi(x)=\psi(x)-\psi(x-e_{i})$, and then $D_{i_{1}^{\varepsilon_{1}},\ldots,i_{k}^{\varepsilon_{k}}}=D_{i_{1}^{\varepsilon_{1}}}\cdots D_{i_{k}^{\varepsilon_{k}}}$ for $i_{1},\ldots,i_{k}\in\{1,\ldots,d\}$ and $\varepsilon_{1},\ldots,\varepsilon_{k}\in\{+1,-1\}$. For instance, we can write the graph Laplacian as $\Delta_{d}=\sum_{i=1}^{d}D_{i^{+},i^{-}}$ We chose to focus on the $k$-th differences defined in (2.6) for simplicity, and refer to Remark 4.1 below for a discussion on how to adapt the proofs (see in particular (4.17)).
The same regularity results also hold for the corresponding continuous first eigenfunction $\varphi_{1}$. This seems to be standard (we point to [kondrat1983boundary] for an overview), but we lack a proper reference in such a simple form. We recover these estimates easily, thanks to our coupling arguments.
Theorem 2.4.
Assume that $\Omega$ satisfies Assumption 1. Then there is a constant $C>0$ (that depends on the domain $\Omega$) such that:
  1. (i)
    For any $x\in\Omega$, $|\varphi_{1}(x)|\leq Cd(x,\partial\Omega)$;
  2. (ii)
    For any $k\geq 1$ and any $i_{1},\ldots,i_{k}\in\{1,\ldots d\}$, we have $|D^{(h)}_{i_{1},\ldots,i_{k}}\varphi_{1}(x)|\leq(Ck)^{k}d(x,\partial\Omega)^{1-k}$ for any $x\in\Omega$ with $d(x,\partial\Omega)\geq 2kh$. As a consequence, $\varphi_{1}$ is $C^{\infty}$ in $\Omega$, with
    $\Big{|}\frac{\partial^{k}}{\partial x_{i_{1}}\cdots\partial x_{i_{k}}}\varphi_{1}(x)\Big{|}\leq(Ck)^{k}d(x,\partial\Omega)^{1-k}\,,\qquad\forall\,x\in\Omega\,.$
2.2 Under the exterior cone condition, Assumption 2
We now state how Proposition 2.1 and Theorem 2.2 are modified when using the exterior cone condition of Assumption 2 rather than Assumption 1. The main difference lies in the use of gambler’s ruin estimates: one needs to use estimates that a random walk or a Brownian motion avoids a cone of angle $\alpha$ rather than a ball; for this, [banuelos1997brownian, deblassie1987] and [DW15] turn out to be crucial. Let us state the analogous of Proposition 2.1 in this context.
Proposition 2.5.
Suppose that $\Omega$ satisfies Assumption 2. Then there is a constant $C>0$ and some constant $p=p(\alpha)\in(0,1]$ defined in (2.10) (or (2.11)) below, such that
$0<\phi_{N}(x)\leq C\,\Big{(}\frac{d(x,\partial\Omega_{N})}{N}\Big{)}^{p}\,,\qquad\,\forall x\in\Omega_{N}\,.$ (2.9)
In particular, we also have $\sup_{x\in\Omega_{N}}|\phi_{N}(x)|\leq C\,\mathrm{diam}(\Omega)$.
The exponent $p=p(\alpha)$ depends on the minimal angle of the exterior cone in Assumption 2, and is related to the probability for a Brownian motion or a random walk to avoid a cone $\mathcal{C}_{\alpha}$ with opening angle $\alpha$ (note that this amounts to staying in a cone of opening angle $\theta=\pi-\alpha$); we refer to [banuelos1997brownian] for the Brownian motion case (see also [burkholder1977exit, deblassie1987]) and [DW15] for the random walk case. The value of $p$ is in fact explicit for a generic cone $K$, see [banuelos1997brownian]: consider the Laplace–Beltrami operator on $\mathbb{S}^{d-1}$ and let $\lambda_{K^{c}}$ be its first Dirichlet eigenvalue in $K^{c}\cap\mathbb{S}^{d-1}$, then we have
$p=\sqrt{\lambda_{K^{c}}+\big{(}\tfrac{d}{2}-1\big{)}^{2}}-\big{(}\tfrac{d}{2}-1\big{)}\,.$ (2.10)
In the case of circular cones, there exist an alternative “explicit” expression for $p$, see [burkholder1977exit, p.192-193] or [deblassie1987, §2]. Let $F(a,b;c;z)=\sum_{n=0}^{\infty}\frac{(a)_{n}(b)_{n}}{(c)_{n}}\frac{z^{n}}{n!}$ be the ordinary hypergeometric function, with the standard notation $(x)_{n}=x(x+1)\cdots(x+n)$ for the rising Pochhammer symbol, and define
$h(p,\alpha)=F\big{(}-p,p+d-2;\tfrac{1}{2}(d-1);\cos^{2}(\tfrac{\alpha}{2})\big{)}\,.$
(In comparison to [burkholder1977exit, deblassie1987] we have set $\alpha=\pi-\theta$.) Then, $p(\alpha)$ is the smallest positive zero of the functionaaaIn [burkholder1977exit, deblassie1987], it is defined as the inverse map of $\alpha(p)=\sup\{\alpha\in(0,\pi),h(p,\alpha)=0\}$, which is equivalent. $p\mapsto h(p,\alpha)$, i.e.
$p(\alpha)=\inf\{p>0,h(p,\alpha)=0\}\,.$ (2.11)
In dimension $d=1$ we have the explicit expression $p(\alpha)=\frac{\pi}{2(\pi-\alpha)}$; on the other hand, in dimension $d\geq 3$ we simply know that $\alpha\mapsto p(\alpha)$ is continuous and strictly increasing from $(0,\pi)$ onto $(0,+\infty)$, with $p(\pi/2)=1$.
Remark 2.4.
We can improve Proposition 2.5 slightly, by letting the exponent $p$ in (2.9) depend on $x$. For $x\in\Omega_{N}$, we let $\partial\Omega_{x}=\{y\in\partial\Omega,d(x,y)=d(x,\partial\Omega)\}$ be the set of projected points of $x$ to $\partial\Omega$, and $\alpha_{x}=\inf_{y\in\partial\Omega_{x}}\alpha_{r}(y)$ be the least opening angle of cones associated to $\partial\Omega_{x}$. Then, we can rewrite (2.9) as $|\phi_{N}(x)|\leq C\,\big{(}\frac{d(x,\partial\Omega_{N})}{N}\big{)}^{p_{x}}$, with $p_{x}=p(\alpha_{x})$ defined in (2.10)-(2.11). We could also adapt the following statements in a similar fashion, but we have chosen to keep the parameter $p=\inf\{p_{x}\}$ for simplicity.
Similarly to the previous section, we are also able to control single differences and higher order differences. As above, we state it into two separate results since we will prove Theorems 2.2-2.6 together, as a first step.
Theorem 2.6.
Suppose that $\Omega$ satisfies Assumption 2. Then there is a constant $C>0$ (that depends only on the domain $\Omega$) such that, for any $z\in\Omega_{N}$
$\big{|}\phi_{N}(x)-\phi_{N}(y)\big{|}\leq C\,\Big{(}\frac{d(z,\partial\Omega_{N})}{N}\Big{)}^{p-1}\,\frac{d(x,y)}{N}\,,\qquad\forall\,x,y\in B\big{(}z,\tfrac{1}{2}d(z,\partial\Omega_{N})\big{)}\,,$ (2.12)
with $p\in(0,1]$ from (2.10) (or (2.11)).
Theorem 2.7.
Suppose that $\Omega$ satisfies Assumption 2. Then there is a constant $C>0$ (that depends only on the domain $\Omega$) such that, for any $k\geq 1$ and any $i_{1},\ldots,i_{k}\in\{1,\ldots,d\}$, for any $x\in\Omega_{N}$ with $d(x,\partial\Omega_{N})\geq 4k$, we have
$\big{|}D_{i_{1},\ldots,i_{k}}\phi_{N}(x)\big{|}\leq\frac{(Ck)^{k}}{N^{k}}\,\,\Big{(}\frac{d(x,\partial\Omega_{N})}{N}\Big{)}^{p-k}\,,$ (2.13)
with $p\in(0,1]$ from (2.10) (or (2.11)).
Again, the same regularity results hold for the corresponding continuous first eigenfunction $\varphi_{1}$. This does not seem completely standard for Lipschitz domains (we were not able to find a reference), but it derives easily from our coupling techniques.
Theorem 2.8.
Assume that $\Omega$ satisfies Assumption 2. There is a constant $C>0$ (that depends on the domain $\Omega$) such that, with $p\in(0,1]$ from (2.10) (or (2.11)):
  1. (i)
    for any $x\in\Omega$, $|\varphi_{1}(x)|\leq Cd(x,\partial\Omega)^{p}$;
  2. (ii)
    for all $k\geq 1$ and any $i_{1},\ldots,i_{k}\in\{1,\ldots,d\}^{k}$, $|D^{(h)}_{i_{1},\ldots,i_{k}}\varphi_{1}(x)\big{|}\leq(Ch)^{k}d(x,\partial\Omega)^{p-k}$ for any $x\in\Omega$ with $d(x,\partial\Omega)\geq 4kh$. As a consequence, for any $x\in\Omega$,
    $\Big{|}\frac{\partial^{k}}{\partial x_{i_{1}}\cdots\partial x_{i_{k}}}\varphi_{1}(x)\Big{|}\leq C^{k}d(x,\partial\Omega)^{p-k}\,,\qquad\forall\,x\in\Omega\,.$
We stress that, contrary to Theorems 2.2-2.3, the bounds in Theorems 2.6-2.7 degrade as one gets closer to the boundary; but they are still very useful. These results directly give Theorem 1.1.
2.3 Convergence of eigenfunctions and some consequences
Recall that $(\mu_{1},\varphi_{1})$ are the principal eigenvalue and $L^{2}$-normalized eigenfunction of the Dirichlet Laplacian on $\Omega$ (see (1.1)). One can then consider the function $\varphi_{N}$ defined on $\Omega_{N}$ by
$\varphi_{N}:x\in\Omega_{N}\mapsto\varphi_{1}\Big{(}\frac{x}{N}\Big{)}\,.$
This section is devoted to estimates on how close $\phi_{N}$ is to $\varphi_{N}$ in $L^{2}(\Omega_{N})$ and $L^{\infty}(\Omega_{N})$.
About the $L^{2}$ and $L^{\infty}$ convergence.
Before we turn to our proofs, let us note that we were not able to find a proper reference showing that having a Lipschitz boundary $\partial\Omega$ (i.e. Assumption 2) is sufficient to ensure the $L^{2}$ or $L^{\infty}$ convergence of the principal discrete eigenfunction to its continuous counterpart. The best result we found was that of [bramble1968effects, Cor. 7.1] (see also [kuttler1970finite]), which show the convergence
$\lim_{N\to\infty}\frac{1}{N^{d}}\sum_{x\in\Omega_{N}}\big{(}\phi_{N}(x)-\varphi_{N}(x)\big{)}^{2}=0\,.$
under the condition that the domain $\Omega$ has a smooth ($\mathcal{C}^{2}$) boundary, together with a rate of convergence. Let us mention that [bramble1968effects] considers a finite-difference operator which is adjusted near the boundary, to obtain a $O(h^{2})$ error term. However, as observed in [bolthausen1994localization, Lemma 2.1] and the discussion that follows, if one adapts the techniques of [bramble1968effects] to our setting, one obtains a $O(h)$ error.
In fact, [bramble1968effects, Thm. 6.1 and Thm. 7.2] states a similar convergence result in dimension $d=2$ if $\Omega$ possesses reentrant corners (with a slower convergence rate). We explain below in Appendix A that the proof carries over to dimension $d\geq 2$, under the uniform exterior cone condition (i.e. Assumption 2). Indeed, gambler’s ruin estimates and bounds on $|D\varphi(x)|$ near the reentrant corners are central to [bramble1968effects, Thm. 6.1], and the present paper provides all the necessary estimates to adapt it.
Theorem 2.9 ($L^{2}$ convergence).
If the domain has a Lipschitz boundary, i.e. if Assumption 2 holds, then we have the $L^{2}$ convergence for the principal eigenfunction: there is a constant $\kappa>0$ such that
$\frac{1}{N^{d}}\sum_{x\in\Omega_{N}}\big{(}\phi_{N}(x)-\varphi_{N}(x)\big{)}^{2}\leq\kappa N^{-p}\,,$
where $p\in(0,1]$ is the exponent from (2.10) (or (2.11)); $p=1$ if Assumption 1 holds.
Bramble and Hubbard also show the sup-norm convergence, and their method can also be adapted: we state the following result.
Theorem 2.10 ($L^{\infty}$ convergence).
Suppose that $\Omega$ satisfies Assumption 2, then
$\lim_{N\to\infty}\sup_{x\in\Omega_{N}}\big{|}\phi_{N}(x)-\varphi_{N}(x)\big{|}=0\,.$
Moreover, if Assumption 1 holds, the convergence is at rate $N^{-1}$.
Remark 2.5.
Of course these results have their counterparts in the notation of the discrete Dirichlet problem (1.2), considering the convergence of $\varphi_{1}^{(h)}$ to $\varphi_{1}$ in $L^{2}(\Omega^{(h)})$ and $L^{\infty}(\Omega^{(h)})$, with $h\vcentcolon=\frac{1}{N}$. Theorems 2.9 and 2.10 translate into
$h^{d}\sum_{x\in\Omega^{(h)}}\big{|}\varphi_{1}(x)-\varphi_{1}^{(h)}(x)\big{|}^{2}\leq\kappa h^{p}\qquad\text{and}\qquad\lim_{h\to 0}\sup_{x\in\Omega^{(h)}}\big{|}\varphi^{(h)}_{1}(x)-\varphi_{1}(x)\big{|}=0\,.$
We refer to Appendix A for a detailed discussion on the proof of [bramble1968effects]: we provide a summary of the method and how our results of Section 2 can be used to obtain Theorems 2.9 and 2.10. Let us mention that rate of convergence for the $L^{\infty}$ convergence in a Lipschitz domain is intricate, so we have prefered to stick to a simpler statement (again, see Appendix A details).
Some consequences in the bulk.
Theorems 2.9-2.10 allow us to derive a few corollaries of the properties of $\phi_{N}$ in the so-called bulk of $\Omega_{N}$. For any $\eta>0$, define
$\Omega_{N}^{\eta}\vcentcolon=\{x\in\Omega_{N},d(x,\partial\Omega_{N})>\eta N\}\,,$
referred to as the bulk of $\Omega_{N}$.
Corollary 2.11.
Assume that $\Omega$ satisfies Assumption 2. Then, for any $\eta>0$ we have
$\sup_{x\in\Omega_{N}^{\eta}}\bigg{|}\frac{\phi_{N}(x)}{\varphi_{N}(x)}-1\bigg{|}\xrightarrow{N\to\infty}0\,.$ (2.14)
Proof.
Inside $\Omega$, since the first eigenfunction $\varphi_{1}$ is continuous and positive, we get that for any $\eta>0$, there is some $\kappa_{\eta}>0$ such that $\varphi_{1}(x)\geq\kappa_{\eta}$ uniformly for $x\in\Omega$ with $d(x,\partial\Omega)>\eta$. Then, from Theorem 2.10, we directly have that for any $x\in\Omega_{N}^{\eta}$,
$\sup_{x\in\Omega_{N}^{\eta}}\bigg{|}\frac{\phi_{N}(x)}{\varphi_{N}(x)}-1\bigg{|}\leq\frac{c}{\kappa_{\eta}}\sup_{x\in\Omega_{N}}\big{|}\phi_{N}(x)-\varphi_{N}(x)\big{|}\xrightarrow{N\to\infty}0\,,$
as desired. ∎
Another corollary is that, in the bulk, $\phi_{N}$ is bounded away from 00 provided $N$ large enough. This is an immediate consequence of the positivity of $\varphi$ in the bulk of $\Omega$ (we have $\inf_{x\in\Omega_{N}^{\eta}}\varphi_{N}(x)\geq\kappa_{\eta}$ for some $\kappa_{\eta}>0$ independent of $N$) and Corollary 2.11.
Corollary 2.12.
Assume that $\Omega$ satisfies Assumption 2. Then, for any $\eta>0$, there is a constant $c_{\eta}>0$ such that for all $N$ marge enough,
$c_{\eta}\leq\inf_{x\in\Omega_{N}^{\eta}}\phi_{N}(x)\leq\sup_{x\in\Omega_{N}^{\eta}}\phi_{N}(x)\leq\frac{1}{c_{\eta}}\,.$ (2.15)
We also give a corollary which controls the ratios $\phi_{N}(x)/\phi_{N}(y)$ in the bulk. Note that these ratios appear when considering the confined random walk, see (1.4); in particular, the ratios give the drift felt by the confined random walk.
Corollary 2.13.
Assume that $\Omega$ satisfies Assumption 2. Then, for any $\eta>0$ there is a constant $c_{\eta}>0$ such that, for all large $N$, and all $x,y\in\Omega_{N}^{\eta}$ we have
$\exp\Big{(}-c_{\eta}\frac{d(x,y)}{N}\Big{)}\leq\frac{\phi_{N}(x)}{\phi_{N}(y)}\leq\exp\Big{(}c_{\eta}\frac{d(x,y)}{N}\Big{)}\,.$ (2.16)
In particular, in the bulk $\Omega_{N}^{\eta}$, the ratios $\phi_{N}(x)/\phi_{N}(y)$ are uniformly bounded away from 00 and $\infty$.
Proof.
First of all, as above, we have that $\inf_{x\in\Omega_{N}^{\eta}}\varphi_{N}(x)\geq\kappa_{\eta}$. Thanks to Corollary 2.11, we also have that $\inf_{x\in\Omega_{N}^{\eta}}\phi_{N}(x)\geq\frac{1}{2}\kappa_{\eta}$ for $N$ large enough. Therefore, using Theorem 2.6, we get that $|\tfrac{\phi_{N}(x)}{\phi_{N}(y)}-1|\leq\frac{C_{\eta}}{\kappa_{\eta}}N^{-1}$ uniformly for $x,y\in\Omega_{N}^{\eta}$ with $x\sim y$ (we can take $z=x$, and notice that $d(z,\partial\Omega_{N})\geq\eta N$). Therefore, we end up with
$e^{-c_{\eta}N^{-1}}\leq 1-\frac{C_{\eta}}{\kappa_{\eta}}N^{-1}\leq\frac{\phi_{N}(x)}{\phi_{N}(y)}\leq 1+\frac{C_{\eta}}{\kappa_{\eta}}N^{-1}\leq e^{c_{\eta}N^{-1}}\,,$
where the first and last inequalities holds for $N$ large enough. We then get the conclusion of the corollary by a telescopic product. ∎
2.4 Some comments
Comparison with the literature.
Let us mention that another classical way of approximating eigenvalue problems is the finite element method. This method also yields the uniform convergence of a discrete problem towards the continuous one, see e.g. [ciarlet1973maximum, schatz1977interior] or [ciarlet2002finite] for an overview, however some major differences arise. First, the discretization is done on regular triangulations which are optimized to get a faster convergence rate. Secondly, the discrete functions are in fact transported back to the continuous setting by interpolation over the triangulation. This allows the study of the eigenvalue problem on this particular class of functions while still keeping the original Laplace operator on $\Omega$.
The finite difference method instead fully discretizes the problem and forgets about the continuous setting. This implies to work with discrete functions as well as a discrete Laplace operator, which is much closer to an analysis using Markov chains such as the random walk. Let us stress that the proof of Bramble and Hubbard [bramble1968effects] relies on an analysis of the Green’s function of the random walk killed at the boundary, as well as a Walsch approximation theorem. In comparison, our paper gives some uniform control of the discrete eigenvector $\phi_{N}$.
The main interest of our results lies in the fact that they give important information on the first eigenvector $\phi_{N}$ that are useful on their own (see below), and also in the relative simplicity of their proofs, that rely on probabilistic ideas that might have applications in other contexts. We should point out that coupling techniques have already been used in the context of spectral problems, see [atarburdzy2004, burdzy2006, burdzykendall2000, chen1998, kendall1989] to cite a few. Let us mention in particular [atarburdzy2004], which establish via a coupling argument the “hot spots conjecture” that the Neumann eigenfunction associated with the second Neumann eigenvalue attain its maximum and minimum at boundary points only. In fact, the process associated with the Neumann Laplacian is the Brownian motion reflected at the boundary of $\Omega$ (as opposed to the Brownian motion killed at the boundary for the Dirichlet Laplacian), and [atarburdzy2004] use a type of mirror coupling that needs to apply also when the Brownian motion hits the boundary. In comparison, we will also use a mirror coupling but in a much more simple way.
Let us conclude by saying that, to the best of our knowledge, coupling techniques have only been used in continuous settings (and not for the Dirichlet Laplacian), but never in the context of discrete approximations, as in the present paper. We hope that our results will highlight some of the probabilistic ideas at hand and prove valuable in other contexts.
Some motivations.
One of our main motivation is to study the random walk confined in $\Omega_{N}$ (or in other words, the quasi-stationary dstribution), whose transition kernel is given by (1.4) and involves rations of the form $\phi_{N}(x)/\phi_{N}(y)$. Another interpretation is that the confined random walk is a random walk in conductances $c_{N}(x,y)=\phi_{N}(x)\phi_{N}(y)$, referred to as tilted random walk, so our regularity results appear crucial in studying fine properties of this process.
For instance, the regularity of $\phi_{N}$ can be used to estimate on which time scale the titled random walk is comparable to the simple random walk, i.e. with constant conductances — or with the random walk with conductances $\tilde{c}_{N}(x,y)=\varphi_{N}(x)\varphi_{N}(y)$, using convergence results. As another example, the second author investigates in [bouchot2024confined] the geometry of the confined random walk (i.e. conditioned to remain forever in $\Omega_{N}$), in the bulk of $\Omega_{N}$, giving a coupling between the confined walk and a tilted random interlacement in conductances. Here again, the regularity of $\phi_{N}$ could be useful to determine on which scales the tilted random interlacements are comparable with standard random interlacements. The results of the present article also appear crucial to study covering times of the confined random walk, which is the object of the upcoming work [bouchot-covering].
3 Some preliminaries and first results
In this section, we introduce some probabilistic objects that will appear in the proof, together with useful estimates. In the following $B(x,R)$ denotes the discrete Euclidean ball centered at $x$ of radius $R$.
3.1 Rough bounds on the first eigenvalue
In the proof of Theorem 2.2, we only need very rough (and easy) bounds on the first eigenvalue $\lambda_{N}$, that we collect in the following lemma.
Lemma 3.1.
Let $\Omega$ be an open and bounded set and $\Omega_{N}$ the connected component of 00 in $(N\Omega)\cap\mathbb{Z}^{d}$. Then there are two constants $c_{\Omega},c_{\Omega}^{\prime}$ that depends only on $\Omega$ such that the principal eigenvalue $\lambda_{N}$ of the transition matrix $P_{N}$ of the simple random walk killed upon exiting $\Omega_{N}$ verifies
$1-\frac{c_{\Omega}}{N^{2}}\leq\lambda_{N}\leq 1-\frac{c_{\Omega}^{\prime}}{N^{2}}\,.$
Equivalently, there are constants $10<\gamma_{\Omega}<\gamma_{\Omega}^{\prime}<$ 1 such that $\gamma_{\Omega}\leq(\lambda_{N})^{N^{2}}\leq\gamma^{\prime}_{\Omega}$ uniformly in $N$.
Proof.
The proof is very simple: we simply use that $\Omega$ contains a ball $B^{\rm int}$ and is contained in a ball $B^{\rm ext}$. Since the principal eigenvalue is monotone in the domain, we obtain that $\lambda_{N}$ is sandwiched between the principal eigenvalues of $B^{\rm ext}_{N}$ and $B^{\rm int}_{N}$.
It simply remains to see that the principal eigenvalue $\lambda_{B_{N}}$ of a ball verifies $1-\frac{c}{N^{2}}\leq\lambda_{B_{N}}\leq 1-\frac{c^{\prime}}{N^{2}}$, which is classical, see [lawlerRandomWalkModern2010, Chap. 6.9] or Remark 3.1 below. ∎
3.2 Random walk and confined random walk, Feynman–Kac representation
We let $(X_{n})_{n\geq 0}$ a simple nearest-neighbor random walk on $\mathbb{Z}^{d}$, whose transition probabilities are $p(x,y)\vcentcolon=\frac{1}{2d}\mathbbm{1}_{\{x\sim y\}}$, and let $\mathcal{F}_{n}=\sigma(X_{0},X_{1},\ldots,X_{n})$ be the associated filtration. We denote by $\mathbf{P}_{x}$ the law of the random walk when starting from $x$. For a set $\Lambda\subset\mathbb{Z}^{d}$, we denote
$H_{\Lambda}\vcentcolon=\inf\{n\geq 0,X_{n}\in\Lambda\}$
the hitting time of $\Lambda$, with $\inf\varnothing=+\infty$.
Then, it is well-known that the first eigenvector $\phi_{N}$ is linked to the survival probability for the walk killed on $\partial\Omega_{N}$, see e.g. [lawlerRandomWalkModern2010, Prop. 6.9.1]: fixing $N$ large enough, for any $x\in\Omega_{N}$ we have
$N^{-2d}\sum_{z\in\Omega_{N}}\phi_{N}(z)\times\phi_{N}(x)=\lim_{t\to+\infty}\lambda_{N}^{-t}\,\mathbf{P}_{x}(H_{\partial\Omega_{N}}>t)\,.$ (3.1)
Let us stress that (3.1) shows in particular that, for $x,y\in\Omega_{N}$, $x\sim y$,
$\lim_{t\to\infty}\mathbf{P}_{x}\big{(}X_{1}=y\mid H_{\partial\Omega_{N}}>t\big{)}=\frac{1}{2d}\frac{\lambda_{N}^{-1}\phi_{N}(y)}{\phi_{N}(x)}=\vcentcolon\tilde{p}_{N}(x,y)\,,$ (3.2)
and justifies the definition (1.4) of the confined random walk, which corresponds to the random walk conditioned to remain (forever) in $\Omega_{N}$.
Let us denote $\tilde{\mathbf{P}}_{x}^{N}$ the law of the confined random walk when started from $x$, i.e. the Markov chain on $\Omega_{N}$ with transition probabilities $\tilde{p}_{N}(x,y)$. We then have a useful relation to compare the simple and confined random walks. Consider a set $\Lambda$ which intersects $\Omega_{N}$, and an event $A\in\mathcal{F}_{H_{\Lambda}}$, i.e. an event that depends on the trajectory of the random walk until it hits $\Lambda$. Then, using the transition kernel $\tilde{p}_{N}(x,y)$ from (1.4) and after telescoping the ratios of the $\phi_{N}$’s, we have
[…]
This is in fact a Feynman–Kac type relation for the first eigenfunction, see e.g. [dynkin2002diffusions, §6.3].
Remark 3.1.
Note that (3.1) shows that, for any $x\in\Omega_{N}$, $\lambda_{N}=\lim_{t\to\infty}\mathbf{P}_{x}(H_{\partial\Omega_{N}}>t)^{1/t}$, which also shows that $(\lambda_{N})^{N^{2}}=\lim_{t\to\infty}\mathbf{P}_{0}(H_{\partial\Omega_{N}}>tN^{2})^{1/t}$. For instance, if $\Omega_{N}$ is a ball of radius $rN$ (say centered at 00), one can easily verify using Markov’s property that, for any $k\in\mathbb{N}$,
$\Big{(}\inf_{x\in\Omega_{N},|x|<\frac{1}{2}rN}\mathbf{P}_{x}\big{(}H_{\partial\Omega_{N}}>N^{2},|X_{N^{2}}|<\tfrac{1}{2}rN\big{)}\Big{)}^{k}\\ \leq\mathbf{P}_{0}(H_{\partial\Omega_{N}}>kN^{2})\leq\Big{(}\sup_{x\in\Omega_{N}}\mathbf{P}_{x}\big{(}H_{\partial\Omega_{N}}>N^{2}\big{)}\Big{)}^{k}\,.$
Therefore, by the invariance principle, we find that there are two constants $\gamma<\gamma^{\prime}<1$ (that depend on $r$) such that $\gamma<(\lambda_{N})^{N^{2}}\leq\gamma^{\prime}$, showing for instance that $1-\frac{c}{N^{2}}\leq\lambda_{N}\leq 1-\frac{c^{\prime}}{N^{2}}$.
Let us mention that we can also write an identity similar to (3.3) for the Brownian motion conditioned to stay forever in $\Omega$, recall (1.5). Let $\mathbf{P}_{x}$ denote the law of a Brownian motion $(X_{s})_{s\geq 0}$ starting from $x$ and $\tilde{\mathbf{P}}_{x}$ its version conditioned to remain forever in $\Omega$. By using the Doob’s $\varphi_{1}$-transform of (1.5), letting $H_{\Lambda}=\inf\{s>0,X_{s}\in\Lambda\}$ we have the following Feynman–Kac relation
$\tilde{\mathbf{P}}_{x}(A)=\frac{1}{\varphi_{1}(x)}\mathbf{E}_{x}\Big{[}e^{\mu_{1}H_{\Lambda}}\varphi_{1}(X_{H_{\Lambda}})\mathbbm{1}_{\{H_{\Lambda}<H_{\partial\Omega}\}}\mathbbm{1}_{A}\Big{]}\,,$ (3.4)
for any event $A\in\mathcal{F}_{H_{\Lambda}}$, analogously to (3.3).
3.3 Gambler’s ruin estimates and a priori bounds on $|\phi_{N}(x)|$
Let us state here a random walk estimate, based on classical gambler’s ruin arguments; its proof is postponed to Section 5.1. We then show how one can deduce Proposition 2.1 from it.
Lemma 3.2.
Under Assumption 1 that $\Omega$ has positive reach, there is a constant $c>0$ such that for all $N$ large enough, all $x\in\Omega_{N}$,
$\mathbf{P}_{x}(H_{\partial\Omega_{N}}>N^{2})\leq C\,\frac{d(x,\partial\Omega_{N})}{N}\,.$
Proof of Proposition 2.1.
Let $t>N^{2}$. Using the Markov property, we can write
$\lambda_{N}^{-t}\mathbf{P}_{x}(H_{\partial\Omega_{N}}>t)\leq\lambda_{N}^{-N^{2}}\mathbf{P}_{x}(H_{\partial\Omega_{N}}>N^{2})\times\lambda_{N}^{-(t-N^{2})}\sup_{z\in\Omega_{N}}\mathbf{P}_{z}(H_{\partial\Omega_{N}}>t-N^{2})\,.$
Notice that using Lemma 3.1, the term $(\lambda_{N})^{-N^{2}}$ is bounded by some universal constant, independent of $N$. Applying Lemma 3.2, we therefore get that
$\lambda_{N}^{-t}\mathbf{P}_{x}(H_{\partial\Omega_{N}}>t)\leq C\,\frac{d(x,\partial\Omega_{N})}{N}\times\sup_{z\in\Omega_{N}}\lambda_{N}^{N^{2}-t}\mathbf{P}^{N}_{z}(H_{\partial\Omega_{N}}>t-N^{2})\,.$
We can now take the limit as $t\to+\infty$ on both sides and exchange the supremum on $z\in\Omega_{N}$ and the limit as $t\to+\infty$ (since $N$ is fixed). Applying (3.1) then yields
$\phi_{N}(x)\leq C\,\frac{d(x,\partial\Omega_{N})}{N}\sup_{z\in\Omega_{N}}\phi_{N}(z)\,.$
Now, it remains to show that there is a constant $C^{\prime}$ such that
$\sup_{z\in\Omega_{N}}\phi_{N}(z)\leq C^{\prime}\,.$ (3.5)
The proof is almost identical to the one of Lemma A.1 in [dingDistributionRandomWalk2021a]: observe that $\phi_{N}(X_{t})\lambda_{N}^{-t}$ is a martingale and use the optimal stopping theorem at time $\tau_{N}\vcentcolon=N^{2}\wedge H_{\partial\Omega_{N}}$ to get
$\begin{split}\phi_{N}(z)&=\mathbf{E}_{z}\big{[}\phi_{N}(X_{\tau_{N}})\lambda_{N}^{-\tau_{N}}\big{]}\\ &=\sum_{w\in\Omega_{N}}\phi_{N}(w)\mathbf{P}_{z}(H_{\partial\Omega_{N}}>N^{2},X_{N^{2}}=w)+\mathbf{E}_{z}\Big{[}0\cdot\lambda_{N}^{-H_{\partial\Omega_{N}}}\mathbbm{1}_{\left\{H_{\partial\Omega_{N}}\leq N^{2}\right\}}\Big{]}\,,\end{split}$
where the second term is zero since $\phi_{N}\equiv 0$ on $\partial\Omega_{N}$. Removing the constraint $H_{\partial\Omega_{N}}>N^{2}$ and using the local limit theorem [lawlerRandomWalkModern2010, Theorem 2.1.3], we get
$\sup_{z\in\Omega_{N}}\phi_{N}(z)\leq\sum_{w\in\Omega_{N}}\phi_{N}(w)\mathbf{P}_{z}(X_{N^{2}}=w)\leq c\sum_{w\in\Omega_{N}}\phi_{N}(w)N^{-d}\,.$
Then, the Cauchy–Schwarz inequality yields
$\sup_{z\in\Omega_{N}}\phi_{N}(z)\leq cN^{-d}\Big{(}|\Omega_{N}|\sum_{w\in\Omega_{N}}\phi_{N}^{2}(w)\Big{)}^{1/2}=c|\Omega_{N}|^{1/2}N^{-d/2}\,,$
where we used the normalization given by (1.3) for the last equality. Since $\Omega$ is bounded, by definition of $\Omega_{N}$ we have that $|\Omega_{N}|^{1/2}N^{-d/2}$ is also bounded uniformly in $N$, hence proving (3.5) and concluding the proof. ∎
Assuming only the uniform cone condition Assumption 2, the main theorem of [DW15] gives the following result, which replaces Lemma 3.2 and somehow shows that it is (strictly) easier for a random walk to avoid a cone than a ball.
Theorem 3.3 ([DW15]).
Under Assumption 2, there is a constant $C>0$ such that, for any $x\in\Omega_{N}$,
$\mathbf{P}_{x}(H_{\partial\Omega_{N}}>N^{2})\leq C\,\Big{(}\frac{d(x,\partial\Omega_{N})}{N}\Big{)}^{p}\,,$
where $p=p(\alpha)\in(0,1]$ is defined in (2.10).
Using this inequality in the above proof, one obtains Proposition 2.5 instead of Proposition 2.1.
Remark 3.2.
Let us stress that the proof of Propositions 2.1 and 2.1 hold in the continuum, i.e. considering a Brownian motion in $\Omega$ instead of the simple random walk in $\Omega_{N}$. Using standard gambler’s ruin estimates for Brownian motion (or [banuelos1997brownian, deblassie1987] under the exterior cone condition), we get the following bounds on the first continuous eigenfunction: for every $x\in\Omega$,
$|\varphi_{1}(x)|\leq Cd(x,\partial\Omega)^{p}\,,$ (3.6)
with $p=1$ under Assumption 1 and $p\in(0,1]$ from (2.10)-(2.11) under Assumption 2. This proves in particular the first part of Theorems 2.4 and 2.8.
4 Regularity of $\phi_{N}$ and $\varphi_{1}$ via couplings
In this section, we prove all of our regularity estimates: we first prove Theorems 2.2 and 2.6; then we explain how the proof works for the Brownian motion and how it adapts to higher-order differences, proving the second part of Theorems 2.4 and 2.8; we then conclude by the proof of the higher-order differences in the discrete case, i.e. Theorems 2.3 and 2.7.
4.1 Single difference estimates and simple mirror coupling
In this section, we prove Theorems 2.2 and 2.6, whose proof we divide into several steps. First of all, we treat the case where $x,y$ are at distance $2$ from each other, which allows us to treat the case with an even distance between $x,y$; we then use it to treat the general case.
When $x,y$ are at distance $2$ from each other.
We work with fixed $x,y\in\Omega_{N}$ at distance $2$ from each other and such that $d(x,\partial\Omega_{N})$ is large enough; in the case where $d(x,\partial\Omega_{N})\leq C$, then one simply uses Proposition 2.1 to get that $|\phi_{N}(x)-\phi_{N}(y)|\leq\max(\phi_{N}(x),\phi_{N}(y))\leq CN^{-(1+d)}$.
Step 1. Rewriting of $|\phi_{N}(x)-\phi_{N}(y)|$. Our starting point is to use the Feynman–Kac relation (3.3) to rewrite $|\phi_{N}(x)-\phi_{N}(y)|$. Let us consider the discrete ball $B(z,R)$ centered at the point $z$ such that $x\sim z$ and $y\sim z$, and of radius
$R\vcentcolon=\frac{d(z,\partial\Omega_{N})}{2}\wedge(\delta_{\Omega}N)\,,$
for some fixed (but small) constant $\delta_{\Omega}$, that only depends on the domain $\Omega$. We also denote $H_{R}\vcentcolon=H_{\partial B(z,R)}$ for simplicity. Then, using the relation (3.3) with $\Lambda=\Omega_{N}\setminus B(z,R)$ so that $H_{\Lambda}=H_{R}\,<+\infty$, we obtain the following Feynman–Kac formula: for any $x\in B(z,R)$
$1=\tilde{\mathbf{P}}^{N}_{x}\Big{(}X_{H_{R}}\in\partial B(z,R)\Big{)}=\frac{1}{\phi_{N}(x)}\mathbf{E}_{x}\Big{[}(\lambda_{N})^{-H_{R}}\,\phi_{N}\big{(}X_{H_{R}}\big{)}\Big{]}\,.$ (4.1)
In particular, for $x,y$ with $x\sim z$, $y\sim z$, we have
$\big{|}\phi_{N}(x)-\phi_{N}(y)\big{|}=\bigg{|}\mathbf{E}_{x}\Big{[}(\lambda_{N})^{-H_{R}}\,\phi_{N}\big{(}X_{H_{R}}\big{)}\Big{]}-\mathbf{E}_{y}\Big{[}(\lambda_{N})^{-H_{R}}\,\phi_{N}\big{(}X_{H_{R}}\big{)}\Big{]}\bigg{|}\,.$ (4.2)
Our goal is now to estimate the difference of expectations in (4.2) thanks to a coupling argument, which works when $x,y$ are at distance $2$ (due to the periodicity of the random walk).
Step 2. Coupling argument. We now construct a coupling of two random walks $X^{(1)}$ and $X^{(2)}$ respectively starting at $x$ and $y$. The coupling that we use is the so-called mirror coupling. The idea is to consider the hyperplane $\mathcal{H}=\mathcal{H}_{x,y}$ which is the mediator between $x$ and $y$ (and goes through $z$). We then take a random walk $X^{(1)}$ that starts from $x$ and we let $X^{(2)}$ be its symmetric with respect to $\mathcal{H}$, which indeed starts from $y$. We let $\tau=\tau_{\mathcal{H}}\vcentcolon=\min\{n,X_{n}^{(1)}=X_{n}^{(2)}\}$ the meeting time of $X^{(1)},X^{(2)}$ (which is also the hitting time of $\mathcal{H}$ for either $X^{(1)}$ or $X^{(2)}$), and we then set $X_{n}^{(1)}=X_{n}^{(2)}$ for all $n\geq\tau$. We denote by $\mathbf{P}_{x,y}$ the joint law of $(X_{n}^{(1)},X_{n}^{(2)})_{n\geq 0}$ that we have just constructed, which ou mirror coupling.
We denote by $H_{R}^{1},H_{R}^{2}$ the hitting times of $\partial B(z,R)$ by $X^{(1)},X^{(2)}$ respectively, and we stress that, on the event $\{\tau\leq\min(H_{R}^{1},H_{R}^{2})\}$ (the coupling is successful), we have that both $X^{(1)},X^{(2)}$ reach $\partial B(z,R)$ at the same time and at the same point; we refer to Figure 1 for an illustration. In fact, under the mirror coupling, we also have $H_{R}^{1}=H_{R}^{2}$, but we keep the index of the walk in the notation for clarity. Therefore, we obtain
[…]
Figure 1: On the left a successful coupling; the two walks exit $B(z,R)$ through the same point. On the right, the coupling fails, and the two walks reach $\partial B(z,R)$ before $\mathcal{H}$ and the coupling fails; the exit points are different (and symmetric).
Now, thanks to Propositions 2.1 and 2.5 and since all points in $B(z,R)$ are at distance $\geq R$ from the boundary, we can also bound
$\phi_{N}(X_{H_{R}})\leq\sup_{w\in\partial B(z,R)}\phi_{N}(w)\leq C\Big{(}\frac{R}{N}\Big{)}^{p}\,,$
with $p=1$ under Assumption 1 and $p\in(0,1]$ from (2.10)-(2.11) under Assumption 2; note that $R\leq d(x,\partial\Omega_{N})$. Putting all together, we get that there is a constant $C^{\prime}>0$ such that, for $x,y$ at distance $2$,
$|\phi_{N}(y)-\phi_{N}(x)|\leq 2C\Big{(}\frac{d(x,\partial\Omega_{N})}{N}\Big{)}^{p}\mathbf{E}_{x}\Big{[}(\lambda_{N})^{-H_{R}}\mathbbm{1}_{\{\tau>H_{R}\}}\Big{]}\,,$ (4.4)
using also the symmetry between $x$ and $y$.
Step 3. Technical estimate and conclusion of the proof. It then only remains to control the expectation appearing in (4.4), which are (standard) simple random walk estimates (by translation and rotation invariance, we can take =$\mathcal{H}=\{z=(z_{1},\ldots,z_{d})\in\mathbb{Z}^{d},z_{1}=0\}$). We postpone the proof of the following lemma to Section 5.3, which collects some other useful random walk estimates.
Lemma 4.1 (Gambler’s ruin).
There is a constant $c_{d}>0$ (that depends only on the dimension) and a constant $C>0$ such that, for any $R\geq 1$ sufficiently large and any $x\in B(0,R/2)$, we have
$\mathbf{E}_{x}\Big{[}e^{c_{d}H_{R}/R^{2}}\mathbbm{1}_{\{\tau_{\mathcal{H}}>H_{R}\}}\Big{]}\leq C\,\frac{d(x,\mathcal{H})}{R}\,,$ (4.5)
where $H_{R}$ is the hitting time of the ball $B(0,R)$ and $\tau_{\mathcal{H}}$ is the hitting time of the hyperplane =$\mathcal{H}=\{z=(z_{1},\ldots,z_{d})\in\mathbb{Z}^{d},z_{1}=0\}$.
Let us stress that the slight difficulty in this lemma comes from the term $e^{c_{d}H_{R}/R^{2}}$, which is unbounded: if this term were absent, it would be a standard gambler’s ruin estimate.
We can then use that $\lambda_{N}^{-1}\leq e^{c_{\Omega}/N^{2}}$ for some constant $c_{\Omega}$, see Lemma 3.1, so that we can bound $(\lambda_{N})^{-H_{R}}\leq e^{c_{\Omega}H_{R}^{x}/N^{2}}\leq e^{c_{\Omega}\delta_{\Omega}^{2}H_{R}/R^{2}}$ in (4.4), recalling also that we considered $R\leq\delta_{\Omega}N$. Therefore, having fixed $\delta_{\Omega}$ small enough (how small depends on the domain $\Omega$), we can apply Lemma 4.1 (with $x=e_{1}$ so $d(x,\mathcal{H})=1$) to obtain that for $x,y$ at distance $2$ from each other,
$|\phi_{N}(y)-\phi_{N}(x)|\leq\frac{C^{\prime}}{R}\,\Big{(}\frac{d(x,\partial\Omega_{N})}{N}\Big{)}^{p}\leq C^{\prime\prime}\;\frac{d(x,\partial\Omega_{N})^{p-1}}{N^{p}}\,.$ (4.6)
Recalling that $p=1$ under Assumption 1 and $p\in(0,1]$ from (2.10)-(2.11) under Assumption 2, this concludes the proof when $x,y$ are at distance $2$ from each other.
When $x,y$ are at an even distance from each other.
We can now easily extend (4.6) to the case where $x,y$ are at an even distance from each other. There exists a constant $C$ such that, for $z\in\Omega_{N}$ with $d(z,\Omega_{N})$ large enough, letting $R=\frac{1}{2}d(z,\partial\Omega_{N})\wedge(\delta_{\Omega}N)$ as above, we have that for any $x,y\in B(z,\frac{1}{2}R)$ at an even distance from each other
$\big{|}\phi_{N}(y)-\phi_{N}(x)\big{|}\leq C\,\frac{d(z,\partial\Omega_{N})^{p-1}}{N^{p}}\,d(x,y)\,.$ (4.7)
Indeed, one simply uses (4.6) together with the triangular inequality, using also that we have $d(w,\partial\Omega_{N})\geq\frac{1}{4}d(z,\partial\Omega_{N})$ for any $w\in B(z,\frac{1}{2}R)$ in a geodesic path from $x$ to $y$ (recall that $p\leq 1$ so we need a lower bound on $d(w,\partial\Omega)$).
The general case.
We now turn to the general case, which cannot directly be dealt with the mirror coupling. The idea is to work with the so-called lazy random walk: fix a lazyness parameter $q\in(0,1/2]$, and consider the $q$-lazy random walk instead of a simple random walk — the idea is to use the lazyness of the random walk to fall back on the case where starting points are at some even distance from each other. Note that the $q$-lazy random walk killed upon exiting $\Omega_{N}$ has transition transition matrix $Q_{N}=qI+(1-q)P_{N}$; in other words, $Q_{N}(x,x)=q$ and $Q_{N}(x,y)=\frac{1-q}{2d}$ for $x\sim y$ with $x,y\in\Omega_{N}$. One can readily see that $Q_{N}\phi_{N}=\lambda_{N}^{(q)}\phi_{N}$ with $\lambda_{N}^{(q)}=q+(1-q)\lambda_{N}$, so that $\phi_{N}$ is again the principal eigenvector of $Q_{N}$, with associated eigenvalue that verifies $1-\lambda_{N}^{(q)}=(1-q)(1-\lambda_{N})$.
Let $z\in\Omega_{N}$ be such that $d(z,\Omega_{N})$ is large enough, and let $R=\frac{1}{2}d(z,\partial\Omega_{N})\wedge(\delta_{\Omega}N)$ as above, and let $x,y\in B(z,\frac{1}{4}R)$. Then, as in (4.2), we obtain that
$\big{|}\phi_{N}(x)-\phi_{N}(y)\big{|}\leq\bigg{|}\mathbf{E}_{x}^{(q)}\Big{[}(\lambda_{N}^{(q)})^{-H_{R}}\,\phi_{N}\big{(}X_{H_{R}}\big{)}\Big{]}-\mathbf{E}_{y}^{(q)}\Big{[}(\lambda_{N}^{(q)})^{-H_{R}}\,\phi_{N}\big{(}X_{H_{R}}\big{)}\Big{]}\bigg{|}\,,$ (4.8)
where here $(X_{n})_{n\geq 0}$ is the $q$-lazy random walk, whose distribution is denoted $\mathbf{P}^{(q)}$.
We now introduce a coupling $\mathbf{P}_{x,y}^{(q)}$ of two $q$-lazy walks $X^{(1)},X^{(2)}$ starting from $x,y$ respectively, in two steps:
  1. (i)
    First, we let $(\xi_{i}^{(1)})_{i\geq 1}$, $(\xi_{i}^{(2)})_{i\geq 1}$ be independent i.i.d. sequences of Bernoulli random variables of parameter $1-q$ and we set $T\vcentcolon=\min\{i\geq 1,\xi_{i}^{(1)}\neq\xi_{i}^{(2)}\}$. Up until time $T$, we set $X_{n}^{(1)}=x+\sum_{i=1}^{n}\xi_{i}^{(1)}U_{i}$ and $X_{n}^{(2)}=y+\sum_{i=1}^{n}\xi_{i}^{(2)}U_{i}$, using the same steps $U_{i}$ (uniform in {±$\{\pm e_{i},1\leq i\leq d\}$) for the two walks; in other words, the two walks evolve in parallel until one stays still ($\xi=0$) but not the other ($\xi=1$).
  2. (ii)
    At time $T$, $X_{T}^{(1)}$ and $X_{T}^{(2)}$ are at an even distance one from the other: we then perform a mirror coupling with respect to the mediator hyperplane $\mathcal{H}=\mathcal{H}_{X_{T}^{(1)},X_{T}^{(2)}}$ between $X_{T}^{(1)},X_{T}^{(2)}$. Note that we might have $X_{T}^{(1)}=X_{T}^{(2)}$, in which case the two walks are coupled starting from time $T$.
Now, we can decompose (4.8) according to whether $T\leq R/8$ or not: we have that
$\big{|}\phi_{N}(x)-\phi_{N}(y)\big{|}\leq\bigg{|}\mathbf{E}_{x,y}^{(q)}\bigg{[}\Big{(}(\lambda_{N}^{(q)})^{-H_{R}^{1}}\,\phi_{N}\big{(}X_{H_{R}^{1}}^{(1)}\big{)}-(\lambda_{N}^{(q)})^{-H_{R}^{2}}\phi_{N}\big{(}X_{H_{R}^{2}}^{(2)}\big{)}\Big{)}\mathbbm{1}_{\{T\leq R/8\}}\bigg{]}\bigg{|}\\ +\mathbf{E}_{x,y}^{(q)}\bigg{[}\Big{|}(\lambda_{N}^{(q)})^{-H_{R}^{1}}\,\phi_{N}\big{(}X_{H_{R}^{1}}^{(1)}\big{)}-(\lambda_{N}^{(q)})^{-H_{R}^{2}}\phi_{N}\big{(}X_{H_{R}^{2}}^{(2)}\big{)}\Big{|}\mathbbm{1}_{\{T>R/8\}}\bigg{]}\,.$ (4.9)
For the first term in (4.9), notice that $T<\min\{H_{R}^{1},H_{R}^{2}\}$ (in fact, $X_{T}^{(1)},X_{T}^{(2)}\in B(z,R/2)$), so that applying the strong Markov property at time $T$, we have that it is equal to
$\bigg{|}\mathbf{E}_{x,y}^{(q)}\bigg{[}(\lambda_{N})^{-T}\mathbbm{1}_{\{T\leq R/8\}}\mathbf{E}_{X_{T}^{(1)},X_{T}^{(2)}}^{(q)}\Big{[}(\lambda_{N}^{(q)})^{-H_{R}^{1}}\,\phi_{N}\big{(}X_{H_{R}^{1}}^{(1)}\big{)}-(\lambda_{N}^{(q)})^{-H_{R}^{2}}\phi_{N}\big{(}X_{H_{R}^{2}}^{(2)}\big{)}\Big{]}\bigg{]}\bigg{|}\\ \leq\mathbf{E}_{x,y}^{(q)}\Big{[}(\lambda_{N})^{-R/8}\big{|}\phi_{N}(X_{T}^{(1)})-\phi_{N}(X_{T}^{(2)})\big{|}\Big{]}\leq 2C\,\frac{d(z,\partial\Omega_{N})^{p-1}}{N^{p}}\,d(x,y)\,.$
For the last inequality, we have used that $(\lambda_{N}^{(q)})^{-R/8}\leq(\lambda_{N}^{(q)})^{-cN^{2}}\leq C^{\prime}$ on one hand and (4.7) on the other hand (together with the fact that $d(X_{T}^{(1)},X_{T}^{(2)})\leq d(x,y)+1\leq 2d(x,y)$, by construction).
For the second term in (4.9), we use Propositions 2.1 and 2.5 to bound $\phi_{N}(X_{H_{R}^{i}}^{(i)})$ by a constant times $R^{p}N^{-p}$, and we apply the Markov property at time $R/8$ to get that it is bounded by a constant times
$\frac{d(z,\partial\Omega_{N})^{p}}{N^{p}}\mathbf{E}_{x,y}^{(q)}\bigg{[}(\lambda_{N})^{-R/8}\mathbbm{1}_{\{T>R/8\}}\Big{(}\mathbf{E}_{X_{R/8}^{(1)}}\big{[}(\lambda_{N}^{(q)})^{-H_{R}}\big{]}+\mathbf{E}_{X_{R/8}^{(2)}}\big{[}(\lambda_{N}^{(q)})^{-H_{R}}\big{]}\Big{)}\bigg{]}\,.$
Note that, as above, $(\lambda_{N})^{-R/8}\leq(\lambda_{N})^{-cN^{2}}$ is bounded by a constant. For the remaining terms, we can use the following lemma, which is classical (we provide a short proof in Section 5.2).
Lemma 4.2 (Exit time).
There is a constant $c_{d}>0$ (that depends only on the dimension) and a constant $C>0$ such that, for any $R\geq 1$ sufficiently large, we have
$\sup_{u\in B(0,R)}\mathbf{E}_{u}\Big{[}e^{c_{d}H_{R}/R^{2}}\Big{]}\leq C\,,$
where $H_{R}$ is the hitting time of the ball $B(0,R)$.
Using that $(\lambda_{N})^{-H_{R}}\leq e^{c_{\Omega}\delta_{\Omega}^{2}H_{R}/R^{2}}$, this lemma shows that $\sup_{u\in B(z,R)}\mathbf{E}_{u}\big{[}\lambda_{N}^{-H_{R}}\big{]}$ is bounded by a constant, provided that $\delta_{\Omega}$ has been fixed small enough. All together, we get that the second term in (4.9) is bounded by a constant times
$\frac{d(z,\partial\Omega_{N})^{p}}{N^{p}}\mathbf{P}_{x,y}^{(q)}\big{(}T>R/8\big{)}\leq C\frac{d(z,\partial\Omega_{N})^{p-1}}{N^{p}}\,,$
where we have used that $T$ is a geometric random variable with parameter $\theta=2q(1-q)<1$, so $\mathbf{P}(T>R/8)=\theta^{R/8}\leq cR^{-1}$. Combining all together, this gives that for any $x,y\in B(z,R/4)$,
$|\phi_{N}(x)-\phi_{N}(y)|\leq C\frac{d(z,\partial\Omega_{N})^{p-1}}{N^{p}}d(x,y)\,,$
as desired. ∎
4.2 Brownian motion and higher order derivatives, multi-mirror coupling
In this section, we adapt the proof to the continuous setting, and we prove the higher order derivative estimates of Theorems 2.4 and 2.8. One simply has to work with the Brownian motion instead of the simple random walk, with actually fewer technical difficulties (in particular, there is no parity issue); in the following, $(X_{s})_{s\geq 0}$ denotes a $d$-dimensional standard Brownian motion, and we use similar notation as for the simple random walk to simplify the exposition.
As a warm-up, let us first explain how to adapt the proof of Theorems 2.2 and 2.6. We then turn to higher-order differences, where the coupling gets a bit more involved.
Single differences, simple mirror coupling.
We let $z\in\Omega$ and set $R=\frac{1}{2}d(z,\partial\Omega)\wedge\delta$, and we consider $H_{R}$ the hitting time of $\partial B(0,R)$ by the Brownian motion $(z+X_{s})_{s\geq 0}$. We start with a similar to (4.1). Using the Doob’s $\varphi_{1}$-transform as in (3.4), we get the following Feynman–Kac formula: for any $x,y\in B(0,\frac{1}{2}R)$,
$|\varphi_{1}(x)-\varphi_{1}(y)|=\bigg{|}\mathbf{E}_{x}\Big{[}e^{\mu_{1}H_{R}}\varphi_{1}(z+X_{H_{R}})\Big{]}-\mathbf{E}_{y}\Big{[}e^{\mu_{1}H_{R}}\varphi_{1}(z+X_{H_{R}})\Big{]}\bigg{|}\,,$ (4.10)
analogously to (4.1). Here, there is no periodicity issue, and we can directly define the mirror coupling: we let $\mathcal{H}$ be the mediator hyperplane between $x$ and $y$ and $X^{(1)}$, $X^{(2)}$ are two mirror Brownian motion with respect to $\mathcal{H}$ until time $\tau\vcentcolon=\inf\{t>0,X_{t}^{(1)}=X_{t}^{(2)}\}$ (which is the hitting time of $\mathcal{H}$ for $X^{(1)}$ and $X^{(2)}$); then $X_{t}^{(1)}=X_{t}^{(2)}$ for any $t\geq\tau$. Under this coupling, whose law is denoted by $\mathbf{P}_{x,y}$, we have that $\varphi_{1}(X_{H_{R}^{1}}^{(1)})=\varphi_{1}(X_{H_{R}^{2}}^{(2)})$ on the event $\{\tau<H_{R}\}$.
Then, analogously to (4.3) using translation and rotation invariance, we obtain that
$|\varphi_{1}(x)-\varphi_{1}(y)|\leq Cd(x,\partial\Omega)^{p}\Big{(}\mathbf{E}_{x}\Big{[}e^{\mu_{1}H_{R}}\mathbbm{1}_{\{\tau>H_{R}\}}\Big{]}+\mathbf{E}_{y}\Big{[}e^{\mu_{1}H_{R}}\mathbbm{1}_{\{\tau>H_{R}\}}\Big{]}\Big{)}\,,$
where we have also used (3.6) to bound $\varphi_{1}(z+X_{H_{R}^{i}}^{(i)})$ for $i=1,2$, uniformly over the exit point.
Now, we use the following result, analogous to Lemmas 4.1-4.2, whose proof is identical (actually simpler) as for the random walk. For any $\mu>0$, there is a constant $C>0$ and some $\delta>0$ such that, for any $R\in(0,\delta]$ and any $x\in B(0,R)$,
$\mathbf{E}_{x}\Big{[}e^{\mu H_{R}}\mathbbm{1}_{\{\tau_{\mathcal{H}}>H_{R}\}}\Big{]}\leq C\frac{d(x,\mathcal{H})}{R}\,,\qquad\text{ and }\qquad\mathbf{E}_{x}\Big{[}e^{\mu H_{R}}\Big{]}\leq C\,,$ (4.11)
where $H_{R}$ is the hitting time of $\partial B(0,R)$ and $\tau_{\mathcal{H}}$ is the hitting time of =$\mathcal{H}=\{0\}\times\mathbb{R}^{d-1}$.
Using the strong Markov property at the hitting time $H_{R/2}^{w}=H_{\partial B(w,R/2)}$ with $w=\frac{x+y}{2}$, we obtain that
$\mathbf{E}_{x}\Big{[}e^{\mu_{1}H_{R}}\mathbbm{1}_{\{\tau>H_{R}\}}\Big{]}\leq\mathbf{E}_{x}\Big{[}e^{\mu_{1}H_{R/2}^{w}}\mathbbm{1}_{\{\tau>H_{R/2}^{w}\}}\mathbf{E}_{X_{H_{R/2}^{w}}}\big{[}e^{\mu_{1}H_{R}}\big{]}\Big{]}\leq C^{2}\,\frac{d(x,y)}{R}\,,$
where we have used (4.11) for the last inequality (using also translation and rotation invariance of the Brownian motion) to bound $\mathbf{E}_{X_{H_{R/2}^{w}}}\big{[}e^{\mu_{1}H_{R}}\big{]}$ by a constant (uniformly on $X_{H_{R/2}^{w}}$), and the fact that $d(x,\mathcal{H})=\frac{1}{2}d(x,y)$. All together, this gives that
$|\varphi_{1}(x)-\varphi_{1}(y)|\leq C\,d(x,\partial\Omega)^{p-1}\,d(x,y)\,,$
which is the continuous analogous of Theorems 2.2 and 2.6. This shows in particular that $D_{i}^{(h)}\varphi_{1}(z)\leq Cd(z,\partial\Omega)^{p-1}$ for any $i\in\{1,\ldots,d\}$, from which one deduces that $|\frac{\partial}{\partial x_{i}}\varphi_{1}(z)|\leq C^{\prime}d(z,\partial\Omega)^{p-1}$, as desired.
Higher-order differences, multi-mirror coupling.
Let $z\in\Omega$ and set $R=\frac{1}{2}d(z,\partial\Omega)\wedge\delta$, as above. Our goal is to show that there exists a constant $C$ (independent of $z$) such that, for any $h$ small enough, for any $i_{1},\ldots,i_{k}\in\{1,\ldots,d\}$, we have that
$D^{(h)}_{i_{1},\ldots,i_{k}}\varphi_{1}(z)\leq(Ck)^{k}d(z,\partial\Omega)^{p-k}\,.$ (4.12)
Letting $h\downarrow 0$ then yields the bounds on $\frac{\partial^{k}}{\partial x_{i_{1}}\cdots\partial x_{i_{k}}}\varphi_{1}$ of Theorems 2.4 and 2.8.
Step 0. Preliminaries. Our starting point is the formula (2.6), which gives that
$(2h)^{k}D^{(h)}_{i_{1},\ldots,i_{k}}\varphi_{1}(z)=\sum_{\alpha\in\{+1,-1\}^{k}}\mathrm{sign}(\alpha)\,\varphi_{1}\big{(}z+x_{\alpha}\big{)}\,,\quad\text{ with }x_{\alpha}=x_{\alpha}^{(h)}=h\sum_{j=1}^{k}\alpha_{j}e_{i_{j}}\,.$
We will work with $h$ small enough so that $kh\leq R/2$, so in particular all points $z+x_{\alpha}$ belong to $B(z,R/2)$. Similarly to (4.10), we therefore can write that
$(2h)^{k}D^{(h)}_{i_{1},\ldots,i_{k}}\varphi_{1}(z)=\sum_{\alpha\in\{+1,-1\}^{k}}\mathrm{sign}(\alpha)\,\mathbf{E}_{x_{\alpha}}\Big{[}e^{\mu_{1}H_{R}}\,\varphi_{1}\big{(}z+X_{H_{R}}\big{)}\Big{]}\,,$
where $H_{R}$ is the hitting time of $\partial B(0,R)$. Our goal is now to construct a coupling $\hat{\mathbf{P}}$ of $2^{k}$ Brownian motions, indexed by $\alpha\in\{-1,+1\}^{k}$ (we denote them $X^{(\alpha)}$), with respective starting points $x_{\alpha}$. With such a coupling, we can rewrite:
$(2h)^{k}D^{(h)}_{i_{1},\ldots,i_{k}}\varphi_{1}(z)=\hat{\mathbf{E}}\bigg{[}\sum_{\alpha\in\{+1,-1\}^{k}}\mathrm{sign}(\alpha)\,e^{\mu_{1}H_{R}^{\alpha}}\,\varphi_{1}\big{(}z+X_{H_{R}}^{(\alpha)}\big{)}\bigg{]}\,,$
where we have denoted $H_{R}^{\alpha}$ the hitting time of $\partial B(0,R)$ by $X^{(\alpha)}$.
Step 1. Properties needed for the coupling. Let us now comment on the desired properties of the relevant coupling. Under $\hat{\mathbf{P}}$, we want to have a coupling time $\tau$ such that, at time $\tau$ there is a partition of $\{+1,-1\}^{k}$ into pairs $(\alpha_{+},\alpha_{-})$ which verify:
  • (i)
    $X^{(\alpha_{+})}_{\tau}=X^{(\alpha_{-})}_{\tau}$;
  • (ii)
    $\mathrm{sign}(\alpha_{+})\mathrm{sign}(\alpha_{-})=-1$.
Coalescing the pairs $(X^{(\alpha_{+})},X^{(\alpha_{-})})$ after time $\tau$, i.e. setting $X^{(\alpha_{+})}_{t}=X^{(\alpha_{-})}_{t}$ for $t>\tau$, we then get that on the event $\tau<\min\{H_{R}^{\alpha_{+}},H_{R}^{\alpha_{-}}\}$ the two Brownian motions $X^{(\alpha_{+})},X^{(\alpha_{-})}$ exit the ball $B(0,R)$ at the same time $H_{R}^{\alpha_{+}}=H_{R}^{\alpha_{-}}$ and at the same point $X_{H_{R}}^{(\alpha_{+})}=X_{H_{R}}^{(\alpha_{-})}$. Therefore, on the event $\tau<\min\{H_{R}^{\alpha_{+}},H_{R}^{\alpha_{-}}\}$, we have that
$\mathrm{sign}(\alpha_{+})\,e^{\mu_{1}H_{R}^{\alpha_{+}}}\,\varphi_{1}\big{(}z+X_{H_{R}}^{(\alpha_{+})}\big{)}+\mathrm{sign}(\alpha_{-})\,e^{\mu_{1}H_{R}^{\alpha_{-}}}\,\varphi_{1}\big{(}z+X_{H_{R}}^{(\alpha_{-})}\big{)}=0\,,$
the two terms cancelling out since they have an opposite sign (recall $\mathrm{sign}(\alpha_{+})\mathrm{sign}(\alpha_{-})=-1$).
All together, with such a coupling, we would get the bound
$(2h)^{k}\big{|}D^{(h)}_{i_{1},\ldots,i_{k}}\varphi_{1}(z)\big{|}\leq\sum_{\alpha\in\{+1,-1\}^{k}}\hat{\mathbf{E}}\Big{[}e^{\mu_{1}H_{R}^{\alpha}}\varphi_{1}\big{(}z+X_{H_{R}}^{(\alpha)}\big{)}\mathbbm{1}_{\{\tau>\min_{\alpha\in\{\pm 1\}^{k}}H_{R}^{\alpha}\}}\Big{]}\\ \leq Cd(z,\partial\Omega)^{p}\sum_{\alpha\in\{+1,-1\}^{k}}\hat{\mathbf{E}}\Big{[}e^{\mu_{1}H_{R}^{\alpha}}\mathbbm{1}_{\{\tau>\min_{\alpha\in\{\pm 1\}^{k}}H_{R}^{\alpha}\}}\Big{]}\,,$
where we have used (3.6) to bound $\varphi_{1}(z+X_{H_{R}}^{(\alpha)})$ uniformly over the exit point $z+X_{H_{R}}^{(\alpha)}$. It then remains to show that our coupling verifies the following property:
[…]
Plugging (4.13) in the display above would yield that $\big{|}D^{(h)}_{i_{1},\ldots,i_{k}}\varphi_{1}(z)\big{|}\leq(Ck)^{k}d(z,\partial\Omega)^{p-k}$, as desired.
Step 2. Construction of the coupling. We let $\hat{\mathbf{P}}$ be the law of $k$ independent standard Brownian motions $\{W^{(j)},1\leq j\leq k\}$ with respective starting points $W^{(j)}_{0}=kh\,e_{i_{j}}$. We then define a generalized mirror coupling (or multi-mirror coupling) as follows.
For any $j$, let us set $W^{(j,+)}=W^{(j)}$ and $W^{(j,-)}$ the mirror image of $W^{(j)}$ with respect to the hyperplane $\mathcal{H}_{j}=\{(x_{1},\ldots x_{d})\in\mathbb{R}^{d},x_{i_{j}}=0\}$; note that $W^{(j,-)}$ starts from $-khe_{i_{j}}$. Then, for $\alpha\in\{-1,+1\}^{k}$, we set
$X^{(\alpha)}\vcentcolon=\frac{1}{k}\sum_{j=1}^{k}W^{(j,\alpha_{j})}\,,$ (4.14)
and we notice that under $\hat{\mathbf{P}}$ the $(X^{(\alpha)})_{\alpha\in\{\pm 1\}^{k}}$ are indeed standard Brownian motions, with starting points $X_{0}^{(\alpha)}=x_{\alpha}=h\sum_{j=1}^{k}\alpha_{j}e_{i_{j}}$.
Define $\tau_{j}=\inf\{t>0,W^{(j)}_{t}\in\mathcal{H}_{j}\}=\inf\{t>0,W^{(j,+)}_{t}=W^{(j,-)}_{t}\}$ the time at which the $j$-th mirror coupling succeed, and let $\tau=\min\{\tau_{j},1\leq j\leq k\}$. Now, for any $1\leq j\leq k$, let us define
$\alpha_{\pm}^{j}=(\alpha_{1},\ldots,\alpha_{j-1},\pm 1,\alpha_{j+1},\ldots,\alpha_{k})\,,$
so $\alpha_{\pm}^{j}$ is simply obtained from $\alpha$ by setting $\alpha_{j}$ to $\pm$. Then for any $j\in\{1,\ldots,k\}$, this provides a ($j$-dependent) partition of $\{-1,+1\}^{k}$ into pairs $\{\alpha_{+}^{j},\alpha_{-}^{j}\}$. Let us stress right away that we have $\mathrm{sign}(\alpha_{+}^{j})\mathrm{sign}(\alpha_{-}^{j})=-1$, since $\alpha_{\pm}^{j}$ only differ by one sign (that of $\alpha_{j}$). Then, in the case where $W^{(j,+)}_{t}=W^{(j,-)}_{t}$, we have that for any $\alpha\in\{-1,+1\}^{k}$
$X^{(\alpha_{+}^{j})}_{t}=\frac{1}{k}\sum_{j^{\prime}\neq j}\alpha_{j}^{\prime}W_{t}^{j,\alpha_{j^{\prime}}}+\frac{1}{k}W^{(j,+)}_{t}=\frac{1}{k}\sum_{j^{\prime}\neq j}\alpha_{j}^{\prime}W_{t}^{j,\alpha_{j^{\prime}}}+\frac{1}{k}W^{(j,-)}_{t}=X^{(\alpha_{-}^{j})}_{t}\,.$
Therefore, at time $\tau$, one can find some $j\in\{1,\ldots,k\}$ (the index of the successful coupling) and an associated partition $\{\alpha_{+}^{j},\alpha_{-}^{j}\}$ which satisfies our requirements (i)-(ii) for the coupling. Note in fact that the partition depends on the index of the mirror coupling which succeed.
Step 3. Control of (4.13). First of all, let us introduce
$T_{R}^{j}=\inf\big{\{}t>0,|W_{t}^{(j)}|=R\big{\}}\quad\text{ and }T\vcentcolon=\min_{1\leq j\leq k}T_{R}^{j}\,.$
Notice that, under the coupling $\hat{\mathbf{P}}$, we have $|X_{t}^{(\alpha)}|\leq\frac{1}{k}\sum_{j=1}^{k}|W_{t}^{(j)}|$ so $|X_{t}^{(\alpha)}|<R$ for all $t<T$. Therefore, we have that $T\leq\min_{\alpha\in\{\pm 1\}^{k}}H_{R}^{\alpha}$, and we can bound the left-hand-side of (4.13) by
[…]
where we have used the strong Markov property at time $T\leq H_{R}^{\alpha}$ for the second inequality. Then, we can use the second inequality in (4.11) with $X^{(\alpha)}$, to get that $\mathbf{E}_{X_{T}^{(\alpha)}}\big{[}e^{\mu_{1}H_{R}^{\alpha}}\big{]}$ is bounded by a constant. For the remaining term, writing that $T\leq\sum_{i=1}^{k}T_{R}^{j}$, we get that
$\hat{\mathbf{E}}\big{[}e^{\mu_{1}T}\mathbbm{1}_{\{\tau>T\}}\big{]}\leq\hat{\mathbf{E}}\Big{[}\prod_{j=1}^{k}e^{\mu_{1}T_{R}^{j}}\mathbbm{1}_{\{\tau_{j}>T_{R}^{j}\}}\Big{]}=\prod_{j=1}^{k}\hat{\mathbf{E}}\Big{[}e^{\mu_{1}T_{R}^{j}}\mathbbm{1}_{\{\tau_{j}>T_{R}^{j}\}}\Big{]}\,,$
using that the $W^{(j)}$ are independent Brownian motions under $\hat{\mathbf{P}}$. Since the starting point of $W^{(j)}$ is $kh\,e_{i_{j}}$, we can apply (4.11) to bound each term in the product by $C\frac{kh}{R}$, which concludes the proof of (4.13) and thus of (4.12). ∎
4.3 Higher order differences in the discrete: Theorems 2.3 and 2.7
Let us stress that one can repeat the same argument as in Section 4.2 in the discrete setting, for the simple random walk. Let us explain how the coupling works and what the differences with Section 4.1 are. In the following, let $z\in\Omega_{N}$ with $d(z,\partial\Omega_{N})\geq 4k$ and denote $R=\frac{1}{4}d(z,\partial\Omega_{N})\wedge\delta$ (large enough).
Consider the following coupling $\hat{\mathbf{P}}$ of $2^{k}$ random walks. Let $(S^{(j)})_{1\leq j\leq k}$ be $k$ independent simple random walks with starting points $S_{0}^{(j)}=e_{i_{j}}$, and let $S^{(j,+)}=S^{(j)}$ and $S^{(j,-)}$ the mirror image of $S^{(j)}$ with respect to the hyperplane $\mathcal{H}_{j}=\{(x_{1},\ldots x_{d})\in\mathbb{Z}^{d},x_{i_{j}}=0\}$. Then, for $\alpha\in\{+1,-1\}^{k}$, we define
$Y^{(\alpha)}=\sum_{j=1}^{k}S^{(j,\alpha_{j})}\,.$
(Note that, contrary to (4.14), we do not divide by $k$, in order to keep $\mathbb{Z}^{d}$-valued random walks.) Now, the starting points of $Y^{(\alpha)}$ are indeed $x_{\alpha}=\sum_{j=1}^{k}\alpha_{j}e_{i_{j}}$, but $(Y_{n}^{(\alpha)})_{n\geq 0}$ are not simple random walks: they still are random walks, but with steps distributed as $V=U_{1}+\cdots+U_{k}$ with $(U_{i})_{1\leq i\leq k}$ simple random walk steps (i.e. independent random variables uniform in {±$\{\pm e_{i},1\leq i\leq d\}$). In other words, we can write $(Y_{n}^{(\alpha)})_{n\geq 0}\stackrel{{\scriptstyle(d)}}{{=}}(X_{kn}^{(\alpha)})_{n\geq 0}$ where $(X_{n}^{(\alpha)})$ is a simple random walk.
This is however not problematic for our purposes: the random walk $(Y_{n})_{n\geq 0}\vcentcolon=(X_{kn})_{n\geq 0}$ have transition matrix $P_{N}^{k}$, so in fact $\phi_{N}$ is still its principal eigenfunction, but with associated eigenfunction $\lambda_{N}^{k}$. Similarly to (4.1), we can therefore write
$\phi_{N}(z)=\mathbf{E}_{z}\Big{[}(\lambda_{N})^{-k\tilde{H}_{R}}\phi_{N}\big{(}z+Y_{\tilde{H}_{R}}\big{)}\Big{]}\,,$
where $\tilde{H}_{R}$ is the exit time of $B(0,R)$ by $z+Y$.
Then, under the coupling $\hat{\mathbf{P}}$, the $k$-th order difference can be rewritten as
$D_{i_{1},\ldots,i_{k}}\phi_{N}(z)=\hat{\mathbf{E}}\bigg{[}\frac{1}{2^{k}}\sum_{\alpha\in\{+1,-1\}^{k}}\mathrm{sign}(\alpha)\,(\lambda_{N})^{-k\tilde{H}_{R}^{\alpha}}\,\phi_{N}\big{(}z+Y_{\tilde{H}_{R}}^{(\alpha)}\big{)}\bigg{]}\,,$ (4.15)
with the obvious notation that $\tilde{H}_{R}^{\alpha}$ is the exit time of $B(0,R)$ by $Y^{(\alpha)}$. As in the continuous setting, letting $\tau_{j}\vcentcolon=\min\{n\geq 0,S^{(j)}_{n}\in\mathcal{H}_{j}\}$ and $\tau=\min\{\tau_{j},1\leq j\leq k\}$, we notice that at the time $\tau$ we can partition {±$\{\pm 1\}^{k}$ into pairs $\{\alpha_{+},\alpha_{-}\}$ with $Y^{(\alpha_{+})}_{\tau}=Y^{(\alpha_{-})}_{\tau}$ and $\mathrm{sign}(\alpha_{+})\mathrm{sign}(\alpha_{-})=-1$, similarly as in the continuous setting. Hence, coalescing the pairs $(Y^{(\alpha_{+})},Y^{(\alpha_{-})})$ after time $\tau$, we get that, on the event $\tau\leq\min_{\alpha\in\{\pm 1\}}\tilde{H}_{R}^{\alpha}$, all the terms in the sum cancel out.
We therefore end up with
$\big{|}D_{i_{1},\ldots,i_{k}}\phi_{N}(z)\big{|}\leq\frac{1}{2^{k}}\sum_{\alpha\in\{+1,-1\}^{k}}\hat{\mathbf{E}}\bigg{[}(\lambda_{N})^{-k\tilde{H}_{R}^{\alpha}}\,\phi_{N}(z+Y_{\tilde{H}_{R}})\mathbbm{1}_{\{\tau>\min_{\alpha\in\{\pm\}^{k}}\{\tilde{H}_{R}^{\alpha}\}\}}\bigg{]}\\ \leq\frac{C}{2^{k}}\,d(z,\partial\Omega_{N})^{p}\,N^{-p}\sum_{\alpha\in\{+1,-1\}^{k}}\hat{\mathbf{E}}\Big{[}(\lambda_{N})^{-k\tilde{H}_{R}^{\alpha}}\mathbbm{1}_{\{\tau>T\}}\Big{]}\,,$
where we have used Proposition 2.1 or 2.5 to bound $\phi_{N}(z+Y_{\tilde{H}_{R}})\leq CR^{p}N^{-p}$ (note that the exit point verifies $z+Y_{\tilde{H}_{R}}\in B(z,2R)$). We have also used that $\tilde{H}_{R}^{\alpha}\geq T\vcentcolon=\min_{1\leq j\leq 1}T^{j}_{R/k}$ where we have defined $T^{j}_{R/k}=\min\{n\geq 0,S_{n}^{(j)}\notin B(0,R/k)\}$. Applying the strong Markov property at time $T\leq\min_{\alpha\in\{\pm 1\}^{k}}\tilde{H}_{R}^{\alpha}$, we get that
$\hat{\mathbf{E}}\Big{[}(\lambda_{N})^{-k\tilde{H}_{R}^{\alpha}}\mathbbm{1}_{\{\tau>T\}}\Big{]}=\hat{\mathbf{E}}\Big{[}(\lambda_{N})^{-kT}\mathbbm{1}_{\{\tau>T\}}\mathbf{E}_{Y_{T}}\big{[}(\lambda_{N})^{-k\tilde{H}_{R}^{\alpha}}\big{]}\Big{]}\leq C\prod_{j=1}^{k}\mathbf{E}\Big{[}(\lambda_{N})^{-kT_{R/k}^{j}}\mathbbm{1}_{\{\tau_{j}>T_{R/k}^{j}\}}\Big{]}\,,$
where we have used Lemma 4.2 to bound the internal expectation by a constant (it easily adapts to the $k$-step random walk) and then the fact that $T\leq\sum_{j=1}^{k}T_{R/k}^{j}$ with the independence of the walks $(S^{(j)})_{1\leq j\leq k}$. Now, we are left with estimates on the simple random walk $S^{(j)}$: thanks to Lemma 4.1 we obtain
[…]
using that $\lambda_{N}^{-k}\leq e^{ck/N^{2}}\leq e^{c_{d}(k/R)^{2}}$ since $R^{2}/N^{2}\leq c_{d}$, provided that $\delta$ is chosen small enough; note that we also used that $R/k\geq 2$.
All together, this proves that $|D_{i_{1},\ldots,i_{k}}\phi_{N}(z)|\leq(Ck)^{k}R^{p-k}N^{-p}$, as desired, concluding the proof of Theorems 2.3 and 2.7. ∎
Remark 4.1 (About directional $k$-th order differences).
In the above, we only dealt with symmetric differences defined in (2.5)-(2.6). As noticed in Remark 2.3, we could also consider directional differences $D_{i^{+}},D_{i^{-}}$. Then, one can obtain a formula analogous to (2.6) for the higher order differences $D_{i_{1}^{\varepsilon_{1}},\ldots,i_{k}^{\varepsilon_{k}}}$, namely
$D_{i_{1}^{\varepsilon_{1}},\ldots,i_{k}^{\varepsilon_{k}}}\psi(x)=\sum_{\alpha\in\{0,1\}^{k}}\mathrm{sign}(\alpha,\varepsilon)\psi\Big{(}x+\sum_{j=1}^{k}\varepsilon_{j}\alpha_{j}e_{i_{j}}\Big{)}\,,$ (4.16)
where $\mathrm{sign}(\alpha,\varepsilon)=(-1)^{m}$ with $m$ the number of $j\in\{1,\ldots,k\}$ such that $\alpha_{j}=0,\varepsilon_{j}=1$ or $\alpha_{j}=1,\varepsilon_{j}=-1$. One could still apply an identity of the type (4.15), but the difference here lies in the fact that the starting points of the different random walks $Y^{(\alpha)}$ should now be $x_{\alpha}=\sum_{j=1}^{k}\varepsilon_{j}\alpha_{j}e_{i_{j}}$, which are not at distance $2$ from each other — hence the mirror coupling of the random walks $(S^{(j)})_{1\leq j\leq k}$ does not fully work.
In order to circumvent this, one needs to work with lazy random walks, as in the last paragraph of Section 4.1. The idea is to construct a coupling of $q$-lazy random walks $(S^{(j,+)},S^{(j,-)})_{1\leq j\leq k}$ which start respectively from $e_{i_{j}}$ and 00, in two steps: first, let the random walks evolve in parallel until they all verify $S^{(j,+)}-S^{(j,-)}=2e_{i_{j}}$ (this takes a geometric number of random walk steps); after this, use the generalized mirror coupling described above. We do not write the details of this coupling and of the proof since it follows from straightforward adaptation of the above and do not bring much insight, but let us state the result that one would obtain. There is a constant $C>0$ such that, for any $k\geq 1$ and any $i_{1},\ldots,i_{k}\in\{1,\ldots,d\}$ and $\varepsilon_{1},\ldots,\varepsilon_{k}\in\{\pm 1\}^{k}$, for any $x\in\Omega_{N}$ with $d(x,\Omega_{N})\geq 4k$, we have
$\big{|}D_{i_{1}^{\varepsilon_{1}},\ldots,i_{k}^{\varepsilon_{k}}}\phi_{N}(x)\big{|}\leq(Ck)^{k}d(x,\partial\Omega_{N})^{p-k}N^{-p}\,.$ (4.17)
In (4.17), we have $p=1$ if Assumption 1 holds or $p\in(0,1]$ as in (2.10)-(2.11) if Assumption 2 holds. This is the analogue of Theorems 2.3 and 2.7 for directional differences.
5 Simple random walk (and Brownian motion) estimates
The main goal of this section is to prove Lemmas 4.1 and 4.2, but we start with the proof of Lemma 3.2, which is a classical gambler’s ruin estimate. We focus on the estimates for simple random walks, since the estimates for the Brownian motion (see (4.11)) are identical (in fact, proofs are simpler). We use the notation $B_{R}\vcentcolon=B(0,R)$ and we will denote $\hat{H}_{R}\vcentcolon=\min\{n\geq 0,X_{n}\in B_{R}\}$ and $\check{H}_{R}\vcentcolon=\min\{n\geq 0,X_{n}\notin B_{R}\}$, or simply $H_{R}=\check{H}_{R}$ if the random walk starts inside $B_{R}$ (according to the previous notation).
5.1 Gambler’s ruin and escaping from large balls
Let us first give a technical lemma on gambler’s ruin probabilities, that we mostly deduce from well-known results (our key reference is [lawler2013intersections]).
Lemma 5.1.
Fix $\alpha>1$. There are constants $c_{1},c_{2}$ (depending on $\alpha-1$), such that for all $R$ large enough, for all $x\in B_{\alpha R}\setminus B_{R}$, we have
$\mathbf{P}_{x}\big{(}\hat{H}_{R}>\check{H}_{\alpha R}\big{)}\leq c_{1}\frac{d(x,B_{R})}{R}\,\,,$ (5.1)
and also
$\mathbf{P}_{x}\big{(}\hat{H}_{R}\wedge\check{H}_{\alpha R}\geq R^{2}\big{)}\leq c_{2}\frac{d(x,B_{R})}{R}\,.$ (5.2)
Remark 5.1.
Lemma 5.1 is useful when the point $x$ is closer to $\partial B_{R}$ than $\partial B_{\alpha R}$. When this is not the case, we can apply the same lemma but with different balls, to obtainbbbOne simply needs to replace $B_{R}$ by a ball $\tilde{B}_{R}$ tangent to $B_{\alpha R}$ such that $d(x,\tilde{B}_{R})=d(x,\partial B_{R})$, and $B_{\alpha R}$ by a ball $\tilde{B}_{(1+\alpha)R}$ with the same center as $\tilde{B}_{R}$ but with a large radius so that $B_{R}\subset\tilde{B}_{(1+\alpha)R}$
$\mathbf{P}_{x}\big{(}\hat{H}_{R}>\check{H}_{\alpha R}\big{)}\leq c_{1}^{\prime}\frac{d(x,\partial B_{\alpha R})}{R}\,,\qquad\mathbf{P}_{x}\big{(}\hat{H}_{R}\wedge\check{H}_{\alpha R}\geq R^{2}\big{)}\leq c_{2}\frac{d(x,\partial B_{\alpha R})}{R}\,.$ (5.3)
Proof of (5.1).
We use Proposition 1.5.10 in [lawler2013intersections], which gives the following estimate in dimension $d\geq 3$: let $x\in B_{\alpha R}\setminus B_{R}$, then
$\mathbf{P}_{x}(\hat{H}_{R}<\check{H}_{\alpha R})=\frac{|x|^{2-d}-\big{(}\alpha R\big{)}^{2-d}+\bar{\mathcal{O}}(R^{1-d})}{R^{2-d}-\big{(}\alpha R\big{)}^{2-d}}\,.$
Injecting $|x|=R+\ell$ with $\ell=|x|-R$ (we may assume that $\frac{\ell}{R}\leq 1/2$ otherwise the bound is trivial), this yields
$\mathbf{P}_{x}(\hat{H}_{R}>\check{H}_{\alpha R})=1-\frac{(1+\ell R^{-1})^{2-d}-\alpha^{2-d}+\bar{\mathcal{O}}(R^{-1})}{1-\alpha^{2-d}}\leq c_{\alpha,d}\big{(}\ell R^{-1}+\bar{\mathcal{O}}(R^{-1})\big{)}\,,$
which is the desired result.
In dimension $d=2$, we have from [lawlerRandomWalkModern2010, Prop. 6.4.1] that, analogously as above,
$\mathbf{P}_{x}(\hat{H}_{R}<\check{H}_{\alpha R})=\frac{\ln(\alpha R)-\ln|x|+\bar{\mathcal{O}}(R^{-1})}{\ln(\alpha R)-\ln R}\,.$
Setting again $|x|=R+\ell$, we get after simplifications that
$\mathbf{P}_{x}(\hat{H}_{R}>\check{H}_{\alpha R})=\frac{\ln(1+\ell R^{-1})+\bar{\mathcal{O}}(R^{-1})}{\ln\alpha}\leq c_{\alpha,d}\big{(}\ell R^{-1}+\bar{\mathcal{O}}(R^{-1})\big{)}\,,$
as needed. ∎
Proof of (5.2).
The proof relies on the usual martingale argument. We fix some $x\in B_{\alpha R}\setminus B_{R}$ such that $|x|-R\leq\delta R$, with $\delta=\delta_{\alpha}<\frac{1}{2}\alpha$ small enough (but fixed) so that in (5.1) we have $\mathbf{P}_{x}(\hat{H}_{R}<\check{H}_{\alpha R})\leq\frac{1}{2}$; note that the bound (5.2) is trivial in the case $|x|-R>(2c_{1})^{-1}R$.
Let us write for simplicity $T\vcentcolon=\min\{\hat{H}_{R},\check{H}_{\alpha R}\}$, and consider the martingale $|X_{t\wedge T}|^{2}-t\wedge T$. Applying the stopping time theorem, we get that
$|x|^{2}=\mathbf{E}_{x}\big{[}|X_{t\wedge T}|^{2}-t\wedge T\big{]}\xrightarrow[]{t\uparrow\infty}\mathbf{E}_{x}\big{[}|X_{T}|^{2}-T\big{]}\,,$ (5.4)
where we have used dominated and monotonous convergence as we took the limit $t\uparrow\infty$. Splitting the last expectation according to whether $X_{T}\in B_{R}$ or not, we have
$|x|^{2}=\mathbf{P}_{x}\big{(}X_{T}\in B_{R}\big{)}\mathbf{E}_{x}\big{[}|X_{T}|^{2}-T\,\big{|}\,X_{T}\in B_{R}\big{]}+\mathbf{P}_{x}\big{(}X_{T}\notin B_{R}\big{)}\mathbf{E}_{x}\big{[}|X_{T}|^{2}-T\,\big{|}\,X_{T}\notin B_{R}\big{]}\,.$
Rearranging the terms, we obtain
$\mathbf{P}_{x}\left(X_{T}\in B_{R}\right)\mathbf{E}_{z}\left[T\,\big{|}\,X_{T}\in B_{R}\right]=\mathbf{P}_{x}\left(X_{T}\in B_{T}\right)\mathbf{E}_{x}\left[|X_{T}|^{2}-|x|^{2}\,\big{|}\,X_{T}\in B_{R}\right]\\ +\mathbf{P}_{x}\left(X_{T}\not\in B_{R}\right)\mathbf{E}_{x}\left[|X_{R}|^{2}-T-|x|^{2}\,\big{|}\,X_{T}\not\in B_{R}\right]\,.$
Since we took $x\notin B_{R}$, on the event $\left\{X_{T}\in B_{R}\right\}$ we have $|X_{T}|^{2}\leq|x|^{2}$, so we can bound the first term by 00. Using also that $T\geq 0$, we end up with
$\mathbf{P}_{x}\left(X_{T}\in B_{T}\right)\mathbf{E}_{x}\left[T\,\big{|}\,X_{T}\in B_{R}\right]\leq\mathbf{P}_{x}\left(X_{T}\notin B_{R}\right)\mathbf{E}_{x}\left[|X_{T}|^{2}-|x|^{2}\,\big{|}\,X_{T}\notin B_{R}\right]\,.$
Note that on the event $\left\{X_{T}\notin B_{R}\right\}$ and since we have fixed $x$ verifying $|x|-R\leq\delta R$, we have $|X_{T}|^{2}-|x|^{2}\leq c_{\alpha}R^{2}$, and recall that we chose $\delta$ small enough so that $\mathbf{P}_{x}\left(X_{T}\in B_{R}\right)\geq\frac{1}{2}$ in (5.1). Therefore, we obtain
$\mathbf{E}_{x}\left[T\,|\,X_{T}\in B_{R}\right]\leq 2c_{\alpha}R^{2}\mathbf{P}_{z}\left(X_{T}\not\in B_{R}\right)\leq 2c_{\alpha}c_{1}R(|x|-R)\,,$ (5.5)
where we have used (5.1) for the last inequality. We thus get
$\begin{split}\mathbf{P}_{x}(T>R^{2})&=\mathbf{P}_{x}(X_{T}\in B_{R})\mathbf{P}_{x}\big{(}T>R^{2}\,|\,X_{T}\in B_{R}\big{)}+\mathbf{P}_{x}(X_{T}\notin B_{R})\mathbf{P}_{x}\big{(}T>R^{2}\,|\,X_{T}\notin B_{R}\big{)}\\ &\leq\mathbf{P}_{x}\big{(}T>R^{2}\,|\,X_{T}\in B_{R}\big{)}+\mathbf{P}_{x}(X_{T}\notin B_{R})\leq 4c_{\alpha}c_{1}\frac{|x|-R}{R}+c_{1}\frac{|x|-R}{R}\,,\end{split}$
where in the last line we have used Markov’s inequality together with (5.5) for the first term, and (5.1) for the second term. This concludes the proof of (5.2). ∎
We are now ready to conclude the proof of Lemma 3.2, thanks to Assumption 1.
Proof of Lemma 3.2.
Using the positive reach Assumption 1, there is some $\varepsilon_{0}>0$ for which, for all $x\in\Omega_{N}$, there exists $z\notin\Omega_{N}$ such that the ball $B(z,\varepsilon_{0}N)$ is in $\Omega_{N}^{c}$ and $d(x,B(z,\varepsilon_{0}N))\leq 2d(x,\partial\Omega_{N})$. Now, let $A$ be large enough, so that $\Omega_{N}\subset B(z,AN)$.
Now, we simply observe that $H_{\partial\Omega_{N}}\leq\min\{H_{B(z,\varepsilon_{0}N)},H_{B(z,AN)^{c}}\}$, so that using Lemma 5.1-(5.2) (with $\alpha\vcentcolon=A/\varepsilon_{0}$), we get that
$\mathbf{P}_{x}(H_{\partial\Omega_{N}}>N^{2})\leq c\frac{B(z,\varepsilon_{0}N)}{\varepsilon_{0}N}\leq c^{\prime}\frac{d(x,\partial\Omega_{N})}{N}\,.$
This concludes the proof of Lemma 3.2. ∎
5.2 Proof of Lemma 4.2
The proof is fairly easy. First of all, notice that if we define $\gamma_{R}\vcentcolon=\sup_{v\in B_{R}}\mathbf{P}_{v}(H_{R}>R^{2})$, then we have that $\gamma_{R}\leq\gamma_{d}$ for some constant $\gamma_{d}<1$, uniformly in $R$ large enough — this is due to the invariance principle, we have $\lim_{R\to\infty}\gamma_{R}=\mathbf{P}_{0}(\sup_{s\in[0,1]}|B_{s}|<1)$ with $B$ a Brownian motion.
Then, applying the Markov property iteratively, we have that $\mathbf{P}_{u}(H_{R}\geq jR^{2})\leq(\gamma_{d})^{j}$ for any integer $j$ (uniformly in $u\in B(0,R)$), so that
$\mathbf{E}_{u}\Big{[}e^{c_{d}H_{R}/R^{2}}\Big{]}\leq\sum_{j\geq 0}e^{c_{d}(j+1)}\mathbf{P}_{u}\big{(}H_{R}/R^{2}\in[j,j+1)\big{)}\leq e^{c_{d}}\sum_{j\geq 0}\big{(}e^{c_{d}}\gamma_{d}\big{)}^{j}\,.$
This is bounded by a constant, taking for instance $c_{d}\vcentcolon=\frac{1}{2}\ln\gamma_{d}^{-1}$ so that $e^{c_{d}}\gamma_{d}=\gamma_{d}^{1/2}<1$. ∎
5.3 Proof of Lemma 4.1
The proof will mostly rely on Lemma 5.1. Decomposing over the value of the integer part of $R^{-2}H_{R}$, we get the bound
$\mathbf{E}_{0}\Big{[}e^{c_{d}H_{R}/R^{2}}\mathbbm{1}_{\{\tau_{\mathcal{H}}>H_{R}\}}\Big{]}\leq e^{c_{d}}\mathbf{P}_{0}\big{(}\tau_{\mathcal{H}}>H_{R}\big{)}+\sum_{k\geq 1}e^{c_{d}(k+1)}\mathbf{P}_{0}\big{(}\tau_{\mathcal{H}}>H_{R}\geq kR^{2}\big{)}\,.$
For any $k\geq 1$, we can use the Markov property at time $R^{2}$ to get that
$\mathbf{P}_{0}\big{(}\tau_{\mathcal{H}}>H_{R}\geq kR^{2}\big{)}\leq\mathbf{P}_{0}\big{(}\tau_{\mathcal{H}}\wedge H_{R}\geq R^{2}\big{)}\sup_{v\in B_{R}}\mathbf{P}_{v}\big{(}H_{R}\geq(k-1)R^{2}\big{)}\\ \leq\mathbf{P}_{0}\big{(}\tau_{\mathcal{H}}\wedge H_{R}\geq R^{2}\big{)}\times(\gamma_{d})^{k-1}\,,$
with $\gamma_{R}\vcentcolon=\sup_{v\in B_{R}}\mathbf{P}_{v}(H_{R}>R^{2})\leq\gamma_{d}<1$ is as in the proof of Lemma 4.2. All together, we obtain that
$\mathbf{E}_{0}\Big{[}e^{c_{d}H_{R}/R^{2}}\mathbbm{1}_{\{\tau_{\mathcal{H}}>H_{R}\}}\Big{]}\leq e^{c_{d}}\mathbf{P}_{0}\big{(}\tau_{\mathcal{H}}>H_{R}\big{)}+\mathbf{P}_{0}\big{(}\tau_{\mathcal{H}}\wedge H_{R}\geq R^{2}\big{)}\sum_{k\geq 1}e^{c_{d}}(e^{c_{d}}\gamma_{d})^{k}\,,$
and the last sum is bounded by a constant choosing for instance $c_{d}\vcentcolon=\frac{1}{2}\ln\gamma_{d}^{-1}$. It therefore remains to estimate the two probabilities in the above display.
In order to be in position to apply Lemma 5.1, we introduce some new sets. Let $z\in\mathbb{Z}^{2}$ be such that the ball $\tilde{B}_{R/2}\vcentcolon=B(z,\frac{1}{2}R)$ of radius $R/2$ is tangent to $\mathcal{H}$ on the other side of 00 (so in particular $d(0,\tilde{B}_{R/2})=1$), and let also $\tilde{B}_{R}\vcentcolon=B(z,R)$ and $\tilde{B}_{2R}\vcentcolon=B(z,2R)$. Then, starting from 00, by construction we have that $H_{R}\geq H_{\partial\tilde{B}_{R}}$ on the event $\tau_{\mathcal{H}}>H_{R}$, and also $\tau_{\mathcal{H}}\leq H_{\tilde{B}_{R/2}}$, so that we get
$\mathbf{P}_{0}\big{(}\tau_{\mathcal{H}}>H_{R}\big{)}\leq\mathbf{P}_{0}\big{(}\tau_{\mathcal{H}}>H_{\partial\tilde{B}_{R}}\big{)}\leq\mathbf{P}_{0}\big{(}H_{\tilde{B}_{R/2}}>H_{\partial\tilde{B}_{R}}\big{)}\leq\frac{c}{R}\,.$
For the last inequality, we have used Lemma 5.1-(5.1) (with $\alpha=2$).
On the other hand, we also have by construction that $H_{\tilde{B}_{R/2}}\wedge H_{\partial\tilde{B}_{2R}}\geq\tau_{\mathcal{H}}\wedge H_{R}$, so by Lemma 5.1-(5.2) (with $\alpha=4$), we obtain
$\mathbf{P}_{0}\big{(}\tau_{\mathcal{H}}\wedge H_{R}\geq R^{2}\big{)}\leq\mathbf{P}_{0}\big{(}H_{\tilde{B}_{R/2}}\wedge H_{\partial\tilde{B}_{2R}}\geq(R/2)^{2}\big{)}\leq\frac{c^{\prime}}{R}\,.$
This concludes the proof of Lemma 4.1. ∎
Acknowledgments. The authors would like to thank Antoine Mouzard for numerous, enthusiastic, enlightening discussions. We would also like to thank Karine Beauchard and Monique Dauge for helpful pointers to some of the literature. Both authors also acknowledge the support of grant ANR Local (ANR-22-CE40-0012).
Appendix A Outlining and adapting Bramble and Hubbard’s proof of $L^{2}$ and $L^{\infty}$ convergence
In this appendix, we prove Theorems 2.9 and 2.10. The proof of Bramble and Hubbard [bramble1968effects] applies with very minor modifications, but we provide a complete summary for the sake of completeness. In order to make the proof closer to that of [bramble1968effects], we work with the notation of the discrete Dirichlet problem (1.2): we consider the discrete Laplacian $\Delta^{(h)}$ on $\Omega^{(h)}$ with Dirichlet boundary conditions, and we denote by $\varphi_{1}^{(h)}$ its principal eigenfunction. Our main goal is therefore to show
$h^{d}\sum_{x\in\Omega^{(h)}}\big{|}\varphi_{1}^{(h)}(x)-\varphi_{1}(x)\big{|}^{2}\leq\kappa h^{p}\,,$ (A.1)
in a first time, and
$\lim_{h\to 0}\sup_{x\in\Omega^{(h)}}\big{|}\varphi^{(h)}_{1}(x)-\varphi_{1}(x)\big{|}=0\,.$ (A.2)
in a second time.
The proof of the $L^{2}$ convergence (A.1) mostly consists in controlling the difference $|\mu_{1}^{(h)}-\mu_{1}|$ of the asociated eigenvalues, from which the procedure described in [bramble1968effects, Section 7] will allows us to conclude. It heavily relies on a key estimate on the derivative of $\varphi_{1}$ near the boundary (see [bramble1968effects, Lem. 6.1-6.2]) that Theorem 2.8 will provide. The other key estimate is a bound on $\varphi_{1}^{(h)}$ near the boundary, which Proposition 2.5 provides. The $L^{\infty}$ convergence will follow from applying almost verbatim the proof of [bramble1968effects, Thmeorem 7.1]. We now provide a detailed outline of the proof.
A.1 Eigenvalue convergence
Before investigating the convergence of discrete eigenvalues, we first recall the minimum-maximum property: letting $\mu_{j}^{(h)}$ the $j$-th discrete eigenvalue, we have
$\mu^{(h)}_{j}=\min_{f_{1},\dots,f_{j}}\max_{\alpha_{1},\dots,\alpha_{j}}\frac{1}{\|f\|_{2,h}^{2}}h^{d-2}\sum_{x\in h\mathbb{Z}^{d}}\sum_{i=1}^{d}\big{(}f(x+he_{i})-f(x)\big{)}^{2}\,,$ (A.3)
where $\alpha_{1},\dots,\alpha_{j}$ are real numbers, $f_{1},\dots,f_{j}$ are linearly independent mesh functions that vanish outside $\Omega^{(h)}$, and $f$ is defined as $f=\sum_{k=1}^{j}\alpha_{k}f_{k}$. Note that the sum in (A.3) in over the entire space and can be interpreted as $\|\nabla f\|_{2,h}^{2}$, thus a Dirichlet energy.
To get a bound on the difference of eigenvalues, we may use Weinberger’s method [weinberger1959error] of comparing finite difference domains. Write $\Omega_{\star}^{(h)}$ for the interior of $\Omega^{(h)}$, so that $\Omega_{\star}^{(h)}\cup\partial\Omega_{\star}^{(h)}=\Omega^{(h)}$. We may consider the discrete Dirichlet problem
$\begin{cases}-\Delta^{(h)}w^{(h)}=\mu_{\star}^{(h)}\,w^{(h)}&\quad\text{ on }\Omega_{\star}^{(h)}\,,\\ \qquad\ \ w^{(h)}=0&\quad\text{ on }\partial\Omega_{\star}^{(h)}\,,\end{cases}$ (A.4)
with $w:\Omega^{(h)}\rightarrow\mathbb{R}$. We write $\mu_{\star,j}^{(h)}$ the ordered eigenvalues of the problem (A.4).
Using (3.5)-(3.6) in [bramble1968effects], which are derived from works of Weinberger and others, there are positive constants $c_{1},c_{2}$ such that for all $j\in\{1,\dots|\Omega^{(h)}|\}$,
$(1-c_{1}h^{2})\mu^{(h)}_{j}\leq\mu_{j}\leq(1+c_{2}h^{2})\mu_{\star,j}^{(h)}\,.$ (A.5)
We now only need to control $|\mu^{(h)}_{j}-\mu_{\star,j}^{(h)}|$ in order to prove that $\mu^{(h)}_{j}\to\mu_{j}$. The proof is straightforward with positive reach assumption (Assumption 1), while in the case of Assumption 2, the proof in [bramble1968effects] remains valid. Let us provide a few details.
A proof that $\mu^{(h)}_{1}\to\mu_{1}$ under positive reach assumption.  Note that in (A.3) if we take $j=1$ and $f=\varphi^{(h)}_{1}\mathbbm{1}_{\Omega^{(h)}_{\star}}$ (this is (5.4) in [weinberger1959error]) we get
$\mu^{(h)}_{1}\leq\mu_{\star,1}^{(h)}\leq\bigg{(}\mu^{(h)}_{1}+h^{d-2}\sum_{x\in\partial\Omega_{\star}^{(h)}}\varphi^{(h)}_{1}(x)^{2}\bigg{)}\,\Bigg{/}\,\bigg{(}1-h^{d}\sum_{z\in\Omega^{(h)}\setminus\Omega_{\star}^{(h)}}\varphi^{(h)}_{1}(z)^{2}\bigg{)}\,.$
Observe that both $|\Omega^{(h)}\setminus\Omega_{\star}^{(h)}|$ and $|\partial\Omega_{\star}^{(h)}|$ are of order $h^{1-d}$. Therefore, using the bound on $\varphi^{(h)}_{1}$ given by Proposition 2.1, this yields
$\mu^{(h)}_{1}\leq\mu_{\star,1}^{(h)}\leq\frac{\mu^{(h)}_{1}+ch}{1-c^{\prime}h^{3}}\leq\mu^{(h)}_{1}+c^{\prime\prime}h\,.$
Recalling (A.5) this gives $(1-c_{1}h^{2})\mu^{(h)}_{1}\leq\mu_{1}\leq(1+c_{2}h^{2})(\mu^{(h)}_{1}+c^{\prime\prime}h)$, hence proving
$|\mu_{1}-\mu^{(h)}_{1}|\leq Kh\,.$ (A.6)
Generalising to other Lipschitz domains or other eigenvalues.  Under Assumption 2, applying the previous strategy yields $\mu^{(h)}_{1}\leq\mu_{\star,1}^{(h)}\leq\mu^{(h)}_{1}+c^{\prime\prime}h^{2p-1}$ for $p=p(\alpha)$ appearing in Proposition 2.5. However, from the discussion that follows Proposition 2.5, the exponent $p$ may be arbitrarily small, which means that $h^{2p-1}$ may diverge as $h\to 0$. Still, in the case of Lipshitz domains, one can prove (see [bramble1968effects, Theorem 4.1]) that $|\mu^{(h)}_{j}-\mu_{\star,j}^{(h)}|\to 0$ using a variation of the previous proof. This in fact only requires a Walsh approximation theorem for the domain $\Omega$, which roughly states that harmonic functions on $\Omega$ are close to harmonic functions on a slightly bigger domain; this is satisfied in the case of Lipschitz domains (Bramble and Hubbard refer to [babuska1966numerical, p. 281]).
A.2 From eigenvalue to $L^{2}$ convergence
After obtaining the convergence of eigenvalues, we can deduce a control on the $L^{2}$ error using the eigenvector decomposition. A simple expansion and the $L^{2}$-normalisation of $\varphi^{(h)}_{1}$ implies
$\|\varphi_{1}-\varphi_{1}^{(h)}\|_{L^{2},h}^{2}=1+\|\varphi_{1}\|_{L^{2},h}^{2}-2\langle\varphi_{1},\varphi_{1}^{(h)}\rangle_{h}\,,$ (A.7)
where $\langle\psi,\psi^{\prime}\rangle_{h}=h^{d}\sum_{x\in\Omega^{(h)}}\psi(x)\psi^{\prime}(x)$ is the scalar product in $L^{2}(\Omega^{(h)})$. Controlling the first derivative of $\varphi_{1}$ thanks to Theorem 2.4 or 2.8, we easily get that $\|\varphi_{1}\|_{L^{2},h}^{2}=1+O(h^{p})$ as $h\to 0$. All that is left is thus to prove that $\langle\varphi_{1},\varphi_{1}^{(h)}\rangle_{h}\to 1$ as $h\downarrow 0$, with a (polynomial) control on the decay rate.
We introduce the Green’s function of the simple random walk on $h\mathbb{Z}^{d}$ killed on the boundary of $\Omega^{(h)}$, that is for $x,y\in\Omega^{(h)}\cup\partial\Omega^{(h)}$:
$G^{(h)}(x,y)=\mathbf{E}_{x}\Big{[}\sum_{k=0}^{\tau^{(h)}}\mathbbm{1}_{\left\{X_{k}=y\right\}}\Big{]}\quad\text{where}\quad\tau^{(h)}\vcentcolon=\inf\big{\{}t\geq 0\,:\,X_{t}\in\partial\Omega^{(h)}\big{\}}\,.$
Note that for $x,y\in\Omega^{(h)}$ we have $G^{(h)}(x,y)=G^{(h)}(y,x)$. It is also well-known that $G^{(h)}$ is an inverse to the Laplace operator, in the sense that it satisfies
$-\Delta^{(h)}_{x}G^{(h)}(x,y)\vcentcolon=\frac{h^{-2}}{2d}\sum_{|e|=1}\big{[}G^{(h)}(x+he,y)-G^{(h)}(x,y)\big{]}=h^{-2}\delta_{x,y}\quad\text{ for }x\in\Omega^{(h)}\,,$
as well as $G^{(h)}(x,y)=\delta_{x,y}$ for $x\in\partial\Omega^{(h)}$. From this, we deduce a crucial tool to study $G^{(h)}$ in the form of a discrete Poisson formula: let $f$ be a real-valued function on $\Omega^{(h)}\cup\partial\Omega^{(h)}$ and $x\in\Omega^{(h)}\cup\partial\Omega^{(h)}$, we have
$f(x)=h^{2}\sum_{y\in\Omega^{(h)}}G^{(h)}(x,y)\big{(}-\Delta^{(h)}f(y)\big{)}+\sum_{y\in\partial\Omega^{(h)}}G(x,y)f(y)\,.$ (A.8)
Combining this with the properties of the eigenfunctions/vectors, [bramble1968effects, Section 5] shows the following crucial identity: for $j\geq 1$,
$(\mu^{(h)}_{j}-\mu)\big{\langle}\varphi_{1},\varphi^{(h)}_{j}\big{\rangle}_{h}=\big{\langle}\Delta\varphi_{1}-\Delta^{(h)}\varphi_{1}\,,\,\varphi^{(h)}_{j}\big{\rangle}_{h}=\mu^{(h)}_{j}\big{\langle}\Phi_{h}\,,\,\varphi^{(h)}_{j}\big{\rangle}_{h}\,,$ (A.9)
where we have defined the function $\Phi_{h}$ by
$\Phi_{h}(x)\vcentcolon=h^{2}\sum_{y\in\Omega^{(h)}}G^{(h)}(x,y)\big{[}\Delta\varphi_{1}(y)-\Delta^{(h)}\varphi_{1}(y)\big{]}\,.$
Now, note that since $(\varphi^{(h)}_{j})_{j\geq 1}$ is an orthonormal basis of $L^{2}(\Omega^{(h)})$, we get
$\Big{|}\|\varphi_{1}\|_{L^{2},h}^{2}-\big{\langle}\varphi_{1},\varphi_{1}^{(h)}\big{\rangle}_{h}^{2}\Big{|}=\sum_{j=2}^{|\Omega^{(h)}|}\big{\langle}\varphi_{1},\varphi^{(h)}_{j}\big{\rangle}_{h}^{2}=\sum_{j=2}^{|\Omega^{(h)}|}\bigg{(}\frac{\mu^{(h)}_{j}}{\mu^{(h)}_{j}-\mu}\bigg{)}^{2}\big{\langle}\Phi_{h},\varphi^{(h)}_{j}\big{\rangle}_{h}^{2}\,,$ (A.10)
where we used (A.9) for the second identity. Note that for any $j\geq 2$, $\mu^{(h)}_{j}\geq\mu^{(h)}_{2}$ (the eigenvalues are ordered) and that since $\mu_{1}$ is simple, the convergence of eigenvalues implies that $\mu^{(h)}_{2}\to\mu_{2}>\mu_{1}$. In particular, provided $h$ small enough, the ratio of eigenvalues in (A.10) is non-increasing in $j\geq 2$, and therefore is bounded by a constant independent of $h$ and $j$. Thus, we have
$\big{|}1-\big{\langle}\varphi_{1},\varphi_{1}^{(h)}\big{\rangle}_{h}^{2}\big{|}\leq\big{|}1-\|\varphi_{1}\|_{L^{2},h}^{2}\big{|}+C\sum_{j\geq 2}\big{\langle}\Phi_{h},\varphi^{(h)}_{j}\big{\rangle}_{h}^{2}\leq C\big{(}h^{p}+\|\Phi_{h}\|_{L^{2},h}^{2}\big{)}\,,$
using again that $1-\|\varphi_{1}\|_{L^{2},h}^{2}=\bar{\mathcal{O}}(h^{p})$ as noticed above and the orthonormality of $(\varphi^{(h)}_{j})_{j\geq 1}$.
All that is left is to prove that $\|\Phi_{h}\|_{2,h}^{2}\leq C^{\prime}h^{p}$. Using a Taylor expansion, we easily see that $\Delta\varphi_{1}(y)-\Delta^{(h)}\varphi_{1}(y)$ regroups all the even orders of the expansion of $\varphi_{1}$ in a $h$-neighborhood of $x$. Stopping the Taylor expansion at the $4$-th order, we get
$\Delta\varphi_{1}(y)-\Delta^{(h)}\varphi_{1}(y)=h^{-2}\sum_{\alpha}\Big{[}\frac{D^{\alpha}\varphi_{!}(y)}{\alpha!}+\varepsilon_{\alpha}(h)\Big{]}\,,$
where the sum ranges over all multi-indices $\alpha=(\alpha_{1},\dots,\alpha_{4})\in\mathbb{N}^{4}$ with $\alpha_{i}\geq 0,\alpha_{1}+\dots+\alpha_{4}=4$, and where $\varepsilon_{\alpha}(h)\to 0$ as $h\downarrow 0$ is an error term. Therefore, Theorems 2.4 & 2.8 imply that
$\quad\big{|}\Delta\varphi_{1}(y)-\Delta^{(h)}\varphi_{1}(y)\big{|}\leq Kh^{2}d(y,\partial\Omega)^{p-4}\,,\qquad\forall y\in\Omega^{(h)}_{\geq 4}\,,$ (A.11)
[…]
$\big{|}\Delta\varphi_{1}(y)-\Delta^{(h)}\varphi_{1}(y)\big{|}\leq(2d+1+\mu_{1})\sup_{z\in\Omega^{(h)}_{<4}}|\varphi_{1}(z)|\leq K^{\prime}h^{p}\,.$ (A.12)
where $\Omega^{(h)}_{<4}\vcentcolon=\Omega^{(h)}\setminus\Omega^{(h)}_{\geq 4}$. Combining (A.11) and (A.12), we get
$|\Phi_{h}(x)|\leq h^{2}\sum_{y\in\Omega^{(h)}_{<4}}G^{(h)}(x,y)Kh^{2}d(y,\partial\Omega)^{p-4}+K^{\prime}h^{p}h^{2}\sum_{y\in\Omega^{(h)}\setminus\Omega^{(h)}_{\geq 4}}G^{(h)}(x,y)\,.$
We now use the following estimates on $G^{(h)}$: there are constants $c_{1},c_{2}>0$ such that
$\sum_{y\in\Omega^{(h)}\setminus\Omega^{(h)}_{\geq 4}}G^{(h)}(x,y)\leq c_{1}\,,\qquad h^{4}\sum_{y\in\Omega^{(h)}_{\geq 4}}G^{(h)}(x,y)d(y,\partial\Omega)^{p-4}\leq c_{2}h^{p}\,.$ (A.13)
These can easily be proven using probabilistic methods. For the first term, we see that the sum is the average number of visits to $\Omega^{(h)}_{<4}$, which is dominated by a geometric random variable, hence the first term is bounded by a constant $c_{1}$. For the second term, we decompose the sum over $d(y,\partial\Omega)\in[2^{k},2^{k+1}]h$, so that the second term is bounded by a constant times
$h^{p}\sum_{k}2^{-(4-p)k}G^{(h)}(x,A^{(h)}_{k})\quad\text{with $A^{(h)}_{k}=\big{\{}y\in\Omega^{(h)}\,:\,d(y,\partial\Omega^{(h)})\in[2^{k},2^{k+1}]h\big{\}}$}\,.$
We then observe that $G^{(h)}(x,A^{(h)}_{k})$ is at most $c2^{2k}$ for some $c>0$, as this is the mean time a random walk stays inside a set with characteristic size $2^{k}$. Therefore, the sum over $k$ is bounded, leaving us with $c_{2}h^{p}$.
Combining all of the above, we therefore get $\sup_{x\in\Omega^{(h)}}|\Phi_{h}(x)|\leq ch^{p}$, which concludes the proof of Theorem 2.9. ∎
A.3 Upgrading to a $L^{\infty}$ convergence
We now turn to the proof of Theorem 2.10, applying the ideas of [bramble1968effects, §7]. We use the Poisson formula (A.8) with $\varphi_{1}^{(h)}$ and $\varphi_{1}$ to get that for any $x\in\Omega^{(h)}$
$\varphi^{(h)}_{1}(x)-\varphi_{1}(x)$ $=h^{2}\sum_{y\in\Omega^{(h)}}G^{(h)}(x,y)\big{[}(-\Delta^{(h)}\varphi_{1}(y))-(-\Delta^{(h)}\varphi_{1}^{(h)}(y))\big{]}$
$=h^{2}\sum_{y\in\Omega^{(h)}}G^{(h)}(x,y)\big{[}\Delta\varphi_{1}(y)-\Delta^{(h)}\varphi_{1}(y)+\Delta^{(h)}\varphi_{1}^{(h)}(y)-\Delta\varphi_{1}(y)\big{]}$
$=\Phi_{h}(x)+h^{2}\sum_{y\in\Omega^{(h)}}G^{(h)}(x,y)\big{[}\mu_{1}^{(h)}\varphi_{1}^{(h)}(y)-\mu_{1}\varphi_{1}(y)\big{]}$
$=\Phi_{h}(x)+h^{2}\sum_{y\in\Omega^{(h)}}G^{(h)}(x,y)\big{[}(\mu_{1}^{(h)}-\mu_{1})\varphi_{1}^{(h)}(y)+\mu_{1}(\varphi_{1}^{(h)}(y)-\varphi_{1}(y))\big{]}\,.$
We can now control the last sum using the eigenvalue and $L^{2}$ estimates derived above.
We first note that according to Theorem 2.8 we have that $\|\varphi^{(h)}\|_{\infty}$ is bounded by a constant, so
$\Big{|}(\mu_{1}^{(h)}-\mu_{1})h^{2}\sum_{y\in\Omega^{(h)}}G^{(h)}(x,y)\varphi_{1}^{(h)}(y)\Big{|}$ $\leq c|\mu_{1}^{(h)}-\mu_{1}|h^{2}\sum_{y\in\Omega^{(h)}}G^{(h)}(x,y)\,.$
Then, we can write $h^{2}\sum_{y\in\Omega^{(h)}}G^{(h)}(x,y)=h^{2}\mathbf{E}_{x}\big{[}\tau^{(h)}\big{]}$ with $\tau^{(h)}$ the exit time of $\Omega^{(h)}$ by the simple random walk on $h\mathbb{Z}^{d}$, which is therefore bounded by a constant uniformly in $x\in\Omega^{(h)}$, by the invariance principle.
On the other hand, using Cauchy–Schwartz inequality, we have
$\Big{|}\mu_{1}h^{2}\sum_{y\in\Omega^{(h)}}G^{(h)}(x,y)\big{(}\varphi_{1}^{(h)}(y)-\varphi_{1}(y)\big{)}\Big{|}\leq\mu_{1}h^{2}\Big{(}\sum_{y\in\Omega^{(h)}}G^{(h)}(x,y)^{2}\Big{)}^{1/2}\cdot\|\varphi_{1}^{(h)}-\varphi_{1}\|_{L^{2},h}\,.$
In dimension $d\geq 3$, since $G^{(h)}(x,y)$ is bounded, the first factor above is bounded by a constant times $h^{2}\mathbf{E}_{x}\big{[}\tau^{(h)}\big{]}^{1/2}$, which is itself bounded by a constant times $h$. In dimension $2$, we instead have $G^{(h)}(x,y)\leq c\log h^{-1}$ (see [lawlerRandomWalkModern2010, Proposition 6.3.2]) which leads to a bound $ch^{2}\mathbf{E}_{x}\big{[}\tau^{(h)}\big{]}^{1/2}\log h^{-1}\leq c^{\prime}h\log h^{-1}$.
Combining the above, we find that for $h$ small enough,
$\sup_{x\in\Omega^{(h)}}\big{|}\varphi^{(h)}_{1}(x)-\varphi_{1}(x)\big{|}\leq\sup_{x\in\Omega^{(h)}}\big{|}\Phi_{h}(x)\big{|}+c_{1}|\mu_{1}^{(h)}-\mu_{1}|+c_{2}h\log(h^{-1})^{\mathbbm{1}_{\left\{d=2\right\}}}\|\varphi_{1}^{(h)}-\varphi_{1}\|_{L^{2},h}\,.$
We have seen in the previous sections that all the terms above go to 00 as $h\downarrow 0$: more precisely, we have that $\sup_{x\in\Omega^{(h)}}\big{|}\Phi_{h}(x)\big{|}\leq Ch^{p}$ and $\|\varphi_{1}^{(h)}-\varphi_{1}\|_{2,h}\leq Ch^{p/2}$. We therefore get that
$\sup_{x\in\Omega^{(h)}}\big{|}\varphi^{(h)}_{1}(x)-\varphi_{1}(x)\big{|}\leq Ch^{p}+c_{1}|\mu_{1}^{(h)}-\mu_{1}|\xrightarrow{h\downarrow 0}0\,.$
In particular, under Assumption 1 we have $p=1$ and $|\mu_{1}^{(h)}-\mu_{1}|\leq Kh$ (recall (A.6)), which gives the announced rate of decay. Under Assumption 2, one needs a better control on $|\mu_{1}^{(h)}-\mu_{1}|$.
[…]
[heading=bibintoc]